### v1

# Generating Multi-Table Time Series EHR from Latent Space with Minimal Preprocessing

Abstract
Electronic Health Records (EHR) are time-series relational databases that record patient interactions and medical events over time, serving as a critical resource for healthcare research and applications. However, privacy concerns and regulatory restrictions limit the sharing and utilization of such sensitive data, necessitating the generation of synthetic EHR datasets. Unlike previous EHR synthesis methods—which typically generate medical records consisting of expert-chosen features (e.g., a few vital signs, structured codes only)—we introduce RawMed, the first framework to synthesize multi-table, time-series EHR data that closely resembles raw EHRs. Using text-based representation and compression techniques, RawMed captures complex structures and temporal dynamics with minimal preprocessing. We also propose a new evaluation framework for multi-table time-series synthetic EHRs, assessing distributional similarity, inter-table relationships, temporal dynamics, and privacy. Validated on two open-source EHR datasets, RawMed outperforms baseline models in fidelity and utility. The code is available at https://github.com/eunbyeol-cho/RawMed.
1 Introduction
The digitalization of medical data has accelerated the adoption of Electronic Health Records (EHR), one of the most significant innovations in modern healthcare. EHRs systematically store various medical events, such as prescriptions and test results, along with corresponding timestamps in a multi-table relational database, encompassing categorical, numerical, and text data. Due to these properties, EHRs serve as a crucial resource in various medical AI studies, including clinical prediction, information retrieval, and question answering Goldstein2017OpportunitiesAC ; Shickel18DeepEHR .
However, since EHRs contain sensitive personal information, they are subject to privacy regulations that hinder data sharing and research applications Benitez2010EvaluatingRR . In response, synthetic data has emerged as a promising solution, offering a privacy-preserving alternative that can still support research and clinical applications yan2022multifaceted ; gonzales2023synthetic . Progress in generative models, especially Generative Adversarial Networks (GANs) gan , has opened up new avenues for synthesizing EHR data. Early efforts primarily generated discrete features or heterogeneous features, often overlooking the time-series aspect Choi2017GeneratingMD ; torfi2020corgancorrelationcapturingconvolutionalgenerative . More recent studies incorporate both time-series and heterogeneous features, gradually approaching a closer reflection of real EHR data emrmgan ; ehrsafe ; flexgenehr ; timediff .
Despite this progress, existing EHR synthesis studies still face the following limitations. First, there is a high dependence on feature selection. Most approaches emrmgan ; ehrsafe ; timediff typically select only a subset of tables and columns of the entire EHR database based on domain knowledge, and generate synthetic data exclusively for the features within those selected columns. [^1] In this paper, a column denotes a raw data field in an EHR table, while a feature refers to a processed data representation derived from one or more columns. See Appendix A.1. While this method simplifies the synthesis process, it limits the usefulness of the generated data. For example, if a new research question or analysis requires a feature that was excluded, it is impossible to build predictive models or perform statistical analyses based on that synthetic EHR data. Morever, prior research fullfeature1 ; fullfeature2 ; genhpf suggests that models incorporating a broader range of features tend to achieve higher predictive accuracy, implying that richer synthetic data offers wider utility and more substantial value across various downstream tasks. Second, many existing approaches rely on complex preprocessing steps. Techniques such as numeric binning, one-hot encoding, term normalization, and aggregation are commonly used, but can unintentionally distort the data or obscure meaningful patterns. For instance, aggregating lab values over time may mask sudden anomalies while binning them into discrete ranges may oversimplify subtle trends, reducing the synthetic data’s fidelity for predictive modeling.
Table 1: Maximum number of features and time steps across datasets for each method. [1]: Number of features (see Appendix A), [2]: Number of time steps, [3]: Uses all columns, [4]: Preserves original values.
Method [1] [2] [3] [4]
emrmgan 98 24 $\times$ $\times$
ehrsafe 90 50 $\times$ $\times$
flexgenehr 5,373 48 $\times$ $\times$
timediff 15 276 $\times$ $\times$
RawMed 333,524 243 $\surd$ $\surd$
To address these challenges, our framework, RawMed, synthesizes raw EHRs, multi-table time-series data that preserve all columns and original values in their database form, as illustrated in Figure 1. To implement this approach, RawMed adopts a text-based method, treating EHR data as text to retain original values without additional transformations (e.g., binning, one-hot encoding, normalization). This not only reduces the risk of data distortion, but also makes our method generalizable to various EHR data with different database schemas. However, as time-series data becomes longer, its textual representation typically grows even longer due to subword tokenization, leading to a significant increase in computational complexity.
To mitigate this, we employ Residual Quantization rqvae to compress textualized time-series data, enabling autoregressive modeling in a compressed latent space. This reduces sequence length and computational complexity, allowing RawMed to handle multi-table EHR datasets with numerous columns. Consequently, our framework generates raw EHRs with extensive features and time steps surpassing most existing methods (Table 1). Our key contributions are summarized as follows:
  1. 1.
    First multi-table and time-Series EHR generation. This study introduces RawMed, the first framework for generating raw multi-table time-series EHRs, preserving all columns and original values, and demonstrating the method’s feasibility with three primary tables.
  2. 2.
    Novel evaluation framework for synthetic raw EHRs. Unlike prior work that synthesizes only a subset of features, RawMed generates all elements of raw EHRs, making the quality assessment more challenging. We propose a comprehensive evaluation framework encompassing low-order statistics, downstream utility, multi-table interactions, time-series fidelity, and privacy protection.
  3. 3.
    Open-source validation and code release. We validated RawMed on two open-source EHR datasets, confirming its effectiveness, and plan to release all source code to support further research in synthetic EHR generation.
2 Related work
2.1 EHR data generation
Early research on EHR synthesis primarily focused on generating single data types (e.g., diagnostic codes) using generative models such as GANs and VAEs. Subsequent studies explored mixed data types or incorporated time-series continuity esteban2017real ; biswal2020eva ; zhang2021synteg ; wang2022promptehrconditionalelectronichealthcare ; Theodorou_2023 . However, relatively few studies have addressed both the temporal dynamics and heterogeneous nature of EHR data simultaneously, a gap that has only recently started to gain attention. For example, EMR-M-GAN emrmgan pioneered a GAN-based approach for mixed-type time-series EHR data, using a Dual-VAE and Coupled Recurrent Network (CRN) to capture inter-feature correlations and temporal dynamics. EHR-Safe ehrsafe also adopted a GAN-based framework designed to handle real-world EHR complexities comprehensively, including missing data, diverse feature types, and both static and time-varying attributes. FLEXGEN-EHR flexgenehr integrated static and time-series data, focusing on missing values via optimal transport. TIMEDIFF timediff introduced a hybrid diffusion model capable of generating both continuous and discrete time-series data concurrently. To the best of our knowledge, no existing studies do not utilize all columns from the original EHR database tables. Instead, they rely on a subset of features deemed important by the researchers, and heavily preprocess (e.g., numerical binning, one-hot encoding, normalization, aggregation) the original EHR data to train the generative model.
2.2 Text-based approaches for tabular data generation
With the rapid progress of NLP, studies have emerged that convert tabular data into textual form for generation purposes great ; tabula . This text-based approach offers several advantages: it minimizes complex preprocessing, preserves original values without information loss, and leverages the pre-trained knowledge of language models. Such benefits are particularly evident for datasets like EHRs, which often follow varying database schemas and require substantial preprocessing. Nevertheless, most existing research on tabular data synthesis focuses on single-table settings great ; tabula , with comparatively few studies addressing relational database realtabformer and no studies synthesizing time-series relational database. Moreover, converting tabular data into text increases the length of each row due to subword tokenization, potentially inflating training and inference costs. To mitigate this, RawMed builds on text-based representations but compresses textualized data, enabling efficient autoregressive modeling in a latent space. This approach reduces computational cost while preserving the fidelity of raw, multi-table time-series EHRs across diverse database schemas.
Figure 1: Conceptual overview of the RawMed pipeline. Left: Real-world EHR data. Center: Data generation process. Right: Resulting synthetic data. The bottom illustrates conventional approach focused on feature selection and domain-specific engineering, while the top illustrates RawMed, which minimizes domain-dependent preprocessing to generate raw EHR-like data, thereby enhancing user flexibility and utility.
3 Method
RawMed is a framework for synthesizing multi-table time-series EHRs, capturing complex patient trajectories with minimal preprocessing. Its architecture supports heterogeneous data type and varying event counts using text-based event representation, with two modules: (1) event-level compression for compact event encoding and (2) temporal inter-event modeling for dynamic temporal relationships. This design enables precise synthesis of large-scale raw EHR datasets.
3.1 Data representation and notation
When admitted to a hospital, patients typically undergo a series of clinical events, such as laboratory tests and medication administrations, which are recorded as individual rows in relational tables (e.g., lab, medication). Each table includes columns such as timestamp indicating when the event occurred, item identifying a specific clinical measurement (e.g., glucose) or a drug name (e.g., propofol), and value (e.g., 95) and uom (unit of measure, e.g., mg/dL) for event details. By ordering these events chronologically, a patient’s clinical trajectory is represented as a time-ordered sequence.
Formally, for a patient $p\in\mathcal{P}$, the event sequence is $S^{p}=[e_{1}^{p},e_{2}^{p},\dots,e_{n^{p}}^{p}]$, where $n^{p}$ is the number of events. Each event $e_{i}^{p}$ consists of a timestamp $t_{i}^{p}$, representing time elapsed since hospital admission, an event type $\epsilon_{i}^{p}$ (e.g., lab), and a set of column-value pairs $a_{i}^{p}=\{(c_{j},v_{j})\mid j=1,\dots,m^{\epsilon_{i}^{p}}\}$, where $c_{j}$ is the column name, $v_{j}$ its textualized value, and $m^{\epsilon_{i}^{p}}$ the number of columns for event type $\epsilon_{i}^{p}$. Although the timestamp is a column in the relational table, we isolate $t_{i}^{p}$ from $a_{i}^{p}$ to highlight its role in temporal ordering.
[…]
3.2 Compressing event representations
To address this, RawMed compress the event text embedding $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, where $L$ is the number of tokens and $F$ is the embedding dimension, into a discrete latent representation $\mathbf{z}_{i}^{p}\in\mathbb{R}^{L_{z}\times F_{z}}$, where $L_{z}$ is the latent sequence length and $F_{z}$ is the latent embedding dimension. We adopt neural network-based compression methods, Vector Quantized Variational AutoEncoders (VQ-VAE) vqvae and Residual Quantization (RQ) rqvae , implemented with 1D convolutional neural networks (CNNs).
Encoder and quantization
The encoder Enc transforms $\mathbf{x}_{i}^{p}$ into $\hat{\mathbf{z}}_{i}^{p}=\texttt{Enc}(\mathbf{x}_{i}^{p})\in\mathbb{R}^{L_{z}\times F_{z}}$, consisting of $L_{z}$ latent vectors. Each latent vector, denoted $\hat{\mathbf{z}}\in\mathbb{R}^{F_{z}}$ for simplicity, is quantized to $\mathbf{z}\in\mathbb{R}^{F_{z}}$ by assigning it to the nearest entry in the codebook $C=\{(k,\texttt{lut}(k))\mid k=1,\dots,K\}$, where $\texttt{lut}(k)\in\mathbb{R}^{F_{z}}$ is the embedding for index $k$ [^2] The function lut refers to a look-up table that maps indices to their corresponding embeddings. The quantization is defined as:
$\texttt{VQ}(\hat{\mathbf{z}};C)=\operatorname*{argmin}_{k\in[K]}\|\hat{\mathbf{z}}-\texttt{lut}(k)\|_{2}^{2},\quad\mathbf{z}=\texttt{lut}\big{(}\texttt{VQ}(\hat{\mathbf{z}};C)\big{)}.$
Thus, the quantized latent representation $\mathbf{z}_{i}^{p}\in\mathbb{R}^{L_{z}\times F_{z}}$ comprises $L_{z}$ number of $\mathbf{z}$ vectors. To enhance the expressiveness of the representation, Residual Quantization (RQ) decomposes each latent vector $\hat{\mathbf{z}}\in\mathbb{R}^{F_{z}}$ from $\hat{\mathbf{z}}_{i}^{p}$ into multiple quantized components. Specifically, RQ represents $\hat{\mathbf{z}}$ as a tuple of indices and composes it as:
$\texttt{RQ}(\hat{\mathbf{z}};C,D)=(k_{1},\dots,k_{D})\in[K]^{D},\quad\mathbf{z}=\sum_{d=1}^{D}\texttt{lut}(k_{d}),$
where $D$ is the quantization depth (i.e., $D$=1 for VQ). The RQ process initializes the residual vector as $\mathbf{r}_{0}=\hat{\mathbf{z}}$. For each depth $d=1,\dots,D$, it quantizes the residual to obtain an index $k_{d}=\texttt{VQ}(\mathbf{r}_{d-1};C)$ and computes the next residual as $\mathbf{r}_{d}=\mathbf{r}_{d-1}-\texttt{lut}(k_{d})$. Partial sums $\mathbf{z}^{(d)}=\sum_{m=1}^{d}\texttt{lut}(k_{m})$ yield increasingly refined approximations, with $\mathbf{z}=\mathbf{z}^{(D)}$.
Decoder and loss function
The decoder Dec reconstructs the original embedding as $\hat{\mathbf{x}}_{i}^{p}=\texttt{Dec}(\mathbf{z}_{i}^{p})$. The model is trained by minimizing a combination of the reconstruction loss $\|\mathbf{x}_{i}^{p}-\hat{\mathbf{x}}_{i}^{p}\|_{2}^{2}$ and the commitment loss. The loss of commitment encourages $\texttt{Enc}(\mathbf{x})$ to remain close to $\mathbf{z}$ used during quantization. Details on architecture and training are provided in Appendices D.1 and E.
3.3 Temporal modeling between events
To model temporal relationships between events $e_{i}^{p}$, we transform the patient’s trajectory $S^{p}$ by replacing each event’s text embedding $\mathbf{x}_{i}^{p}$ with its compressed representation $\mathbf{z}_{i}^{p}$. Each $\mathbf{z}_{i}^{p}$ is mapped to a sequence of $L_{z}\cdot D$ discrete indices $k_{i}^{p}\in{[K]}^{L_{z}\times D}$, forming the compressed trajectory $S^{p}_{\text{quantized}}=[(t_{i}^{p},k_{i}^{p})\mid i=1,\dots,n^{p}]$, where $t_{i}^{p}$ is the timestamp of the $i$-th event.
Time tokenization
Each timestamp $t_{i}^{p}$ is tokenized into a fixed-length sequence of $L_{t}$ digits, denoted $\tau_{i}^{p}\in\{0,1,\dots,9\}^{L_{t}}$. With a 10-minute resolution, $t_{i}^{p}$ is divided by 10 and decomposed into its tens, hundreds, and (if needed) thousands digits (e.g., 720 minutes becomes $[7,2]$). This provides explicit temporal context alongside event content.
Sequence representation and autoregressive modeling
Each event $(\tau_{i}^{p},k_{i}^{p})$ is represented as a block of tokens, combining the $L_{t}$ time tokens from $\tau_{i}^{p}$ with the $L_{z}\cdot D$ flattened indices from $k_{i}^{p}$. Concatenating these blocks yields a unified sequence:
$S^{p}_{\text{quantized}}=[\tau_{1}^{p},k_{1}^{p},\tau_{2}^{p},k_{2}^{p},\dots,\tau_{n^{p}}^{p},k_{n^{p}}^{p}],$
with a total length of $n^{p}\times(L_{t}+L_{z}\cdot D)$. This sequence integrates temporal and event-based information, enabling modeling temporal dependencies effectively. We use a Transformer-based model, TempoTransformer, to autoregressively predict each token based on its predecessors. The training objective minimizes the negative log-likelihood:
$\mathcal{L}_{\text{AR}}=-\sum_{p\in\mathcal{P}}\sum_{i=1}^{|S^{p}_{\text{quantized}}|}\log P(s_{i}^{p}\mid s_{1}^{p},\dots,s_{i-1}^{p}),$
where $s_{i}^{p}$ denotes the $i$-th token in $S^{p}_{\text{quantized}}$, and $P(s_{i}^{p}\mid\cdot)$ is the probability of predicting token $s_{i}^{p}$ given all preceding tokens.
3.4 Data sampling and postprocessing
Synthetic patient trajectories are generated by autoregressively sampling token sequences using the trained TempoTransformer with Top-$k$ sampling. Each synthetic sequence $\tilde{S}^{p}$ interleaves time tokens $\tau_{i}^{p}$ and event tokens $k_{i}^{p}$. Sampling is constrained to ensure structural integrity: the first $L_{t}$ tokens per block, representing time, are sampled from the time token vocabulary $[0,9]$, while the subsequent $L_{z}\cdot D$ event tokens are sampled from the event token vocabulary $[K]$, with invalid token probabilities masked. Next, the event tokens are mapped to latent representations $\tilde{\mathbf{z}}_{i}^{p}$ via codebook $C$, decoded by Dec into text embeddings $\tilde{\mathbf{x}}_{i}^{p}$, and converted to text (e.g., lab item Glucose value 95 uom mg/dL). Time tokens are detokenized into timestamps (e.g., $[7,2]\to 720~{}\text{minutes}$). These text-based events and timestamps are then converted to relational tables.
Although the generated data typically reflect the original table and column names, we observed rare but noticeable issues such as misspelled table or column names or extraneous characters in numeric fields. We hypothesize these errors result from both the quantization step and the autoregressive token prediction step. To ensure accurate table construction, we apply the following postprocessing steps: (1) Event-level verification: Each event $e_{i}^{p}$ is validated to ensure that it begins with a valid table name and adheres to the column-value pair format. Misspelled column names are corrected by matching to the closest valid name using Levenshtein distance, and extraneous characters are removed from numeric fields. Only events that pass all checks after corrections are retained for further processing. (2) Patient-level validation: Subsequently, any sequence $S^{p}$ containing any invalid events $e_{i}^{p}$ are discarded. Retained sequences are checked for temporal consistency, removing events starting from any timestamps that do not follow chronological order or fall outside the observation window. Valid sequences are then converted into relational tables, followed by further postprocessing to enforce column-specific constraints (e.g., numeric ranges). These steps collectively ensure that the final tables are structurally and semantically consistent with the expected format. Further details are provided in Appendix D.2, and the postprocessing algorithm is outlined in Algorithm 1.
4 Experiments
4.1 Setup
Dataset
In this study, we used two publicly available EHR datasets, MIMIC-IV mimiciv and eICU eicu . Adhering to the GenHPF genhpf framework, which utilizes all EHR features for diverse prediction tasks, we focused on patients aged 18 years or older who had been admitted to the ICU for more than 12 hours. Based on this framework, we integrated three time-series tables (laboratory tests, prescriptions, and input events) from both datasets. To assess the generalizability of our approach across diverse clinical settings, we conducted additional experiments on the MIMIC-IV-ED mimicived emergency department dataset, with results reported in Appendix F.3. After preprocessing, the datasets comprised 62,704 patients (about 7 million rows) from MIMIC-IV and 143,812 (about 7 million rows) from eICU, with summary statistics detailed in Table 7. The dataset was split into a 9:1 ratio, allocating 90% for training and 10% for testing. The generative model was trained exclusively with the training data. We aimed to generate data across all columns to preserve the original dataset as comprehensively as possible; however, we excluded columns offering no analytical value (e.g., subject_id, orderid). Details on exclusions are provided in Appendix B.3.
Baseline
This study is the first to synthesize raw EHR data, and therefore does not share an identical problem formulation with prior EHR synthesis research, making direct comparisons challenging. Because existing methods rely on predetermined feature selection and complex preprocessing steps, it is infeasible to extend them to the broader generative scope pursued in this work (i.e., all elements in the original database structure). To provide a baseline, we adapt RealTabFormer realtabformer , designed for text-based relational databases rather than time-series relational databases. Built on GPT-2, RealTabFormer is constrained by limited context length, restricting the amount of data it can process. To address this, we concatenate time-series data in chronological order and apply QLoRA dettmers2024qlora to Llama 3.1dubey2024llama , rather than GPT-2 gpt2 , allowing for longer sequences. Moreover, RealTabFormer does not compress data but instead models it directly, effectively serving as a no-compression baseline.
Additionally, we include baselines using SDV sdv (Gaussian Copula-based), RC-TGAN rctgan (GAN-based), and ClavaDDPM clavaddpm (diffusion-based) for multi-table generation. These methods, while not designed for time-series multi-table data, can generate synthetic data by modeling timestamps as standard columns within each table, as is typical in EHR database structures. However, these approaches might struggle to capture the temporal dynamics due to their lack of explicit temporal modeling. Appendix E.1 details the implementation and technical specifications.
4.2 Evaluation framework for synthetic multi-table time-Series EHRs
4.2.1 Existing evaluation frameworks
Evaluation of synthetic multi-table time-series data, particularly for raw EHRs, remains underexplored. Existing frameworks, such as SDMetrics [^3] https://docs.sdv.dev/sdmetrics and Synthcity synthcity , primarily focus on single-table evaluation. In multi-table scenarios, these frameworks typically average metrics across individual tables or merge tables into a single one, obscuring inter-table relationships and temporal dynamics critical to EHR data. To our knowledge, no framework evaluates all components of raw multi-table EHR data while preserving its time-series structure. To address this, we propose an evaluation framework for RawMed that integrates standard metrics (e.g., CDE), adapts metrics to handle the heterogeneity of raw EHRs (e.g., I-CDE), and introduces novel temporal fidelity metrics (e.g., Time Gap and Next Event Prediction). Detailed metric formulations are provided in the Appendix C.
4.2.2 Single-table evaluation
Single-table evaluation assesses the fidelity of synthetic data within individual tables, treating each row as an instance. To capture distributional similarities, we employ both low-order statistics and high-order metrics.
Column-wise Density Estimation (CDE) evaluates the similarity of marginal distributions for each column. For numeric columns, we use the Kolmogorov-Smirnov (KS) statistic (range: [0,1]), and for categorical columns, we use the Jensen-Shannon (JS) distance (range: [0,1]), both with lower values indicating higher similarity. In EHR data, a single column (e.g., amount column) often contains measurements for diverse items (e.g., about 2,000 drugs), identified by the item column (e.g., itemid or drug). To address this, Item-specific Column-wise Density Estimation (I-CDE) filters data for specific items (e.g., creatinine), computes CDE per item, and averages the results across items to ensure the synthetic data preserves the unique characteristics of each clinical entity for precise item-level fidelity.
Pairwise Column Correlation (PCC) and Item-specific Pairwise Column Correlation (I-PCC) assess inter-column dependencies in synthetic data by comparing correlation matrices of real and synthetic tables. For PCC, matrices are computed using Pearson’s correlation coefficient (range: $[-1,1]$) for numeric-numeric pairs, Theil’s U statistic (range: $[0,1]$) for categorical-categorical pairs, and correlation ratio (range: $[0,1]$) for categorical-numeric pairs. The mean absolute difference ($\mu_{\mathrm{abs}}$) quantifies dependency fidelity by averaging element-wise matrix differences. I-PCC extends this by filtering data for each item, computing $\mu_{\mathrm{abs}}$ per item’s correlation matrix, and averaging across items for precise item-level fidelity.
For high-order dependencies, Predictive Similarity evaluates the ability of synthetic data to model complex, non-linear dependencies in single-table data through predictive performance. An XGBoost model is trained with each column as the target and remaining columns as inputs, with performance evaluated on real test data. For numeric targets, we use Symmetric Mean Absolute Percentage Error (SMAPE; range: [0,200]); for categorical targets, we use classification error rate (ER; range: [0,100]). A smaller performance gap between models trained on synthetic data compared to those trained on real data indicates effective capture of high-order dependencies.
4.2.3 Time-series multi-table evaluation
EHR data are typically stored across multiple tables, each containing time-series events for individual patients. To address the complex interactions among these tables, we preserve both the original structure and time-series aspects. Each patient identified by a primary key (e.g., stay_id), is treated as an instance.
We evaluate the Clinical Utility of synthetic data for downstream predictive tasks in a multi-table EHR setting. Unlike single-table data, multi-table EHR data exhibit significant variability in event types and lengths per patient, necessitating comprehensive representations. To address this variability, we employ two methods: GenHPF genhpf , which lists all events sequentially as a single textual sequence, and MEDS-TAB medstab , which aggregates events into fixed time intervals (see Appendix E.2.1 for details). For evaluation, we define 11 clinical prediction tasks (see Table 10), and report the Area Under the Receiver Operating Characteristic Curve (AUROC) for models trained on synthetic data compared to those trained on real data, using real test data.
We also perform Membership Inference Attacks (MIA) mia to evaluate privacy leakage in synthetic data by measuring distances between synthetic samples and the training/test dataset to infer membership, where successful identification indicates inadequate privacy protection.
To evaluate the temporal fidelity, we introduce two metrics: Time Gap and Event Count. Time Gap quantifies the similarity of distributions of time gaps between consecutive events for each patient, employing the KS statistic (range: [0,1]) on timestamps. Event Count assesses the distribution of event counts per patient, also using the KS statistic, to evaluate whether synthetic data accurately reflects the frequency of clinical events.
Finally, Next Event Prediction evaluates temporal sequence dynamics using an LSTM lstm -based model to predict the next event’s item or drug name, formulated as a multi-label classification task for concurrent events. The model is trained separately on synthetic and real data, with performance evaluated on real test data using the F1 score, quantifying the synthetic data’s ability to capture event sequence patterns.
Table 2: Results of single-table evaluation on MIMIC-IV and eICU datasets. Metrics are averaged across columns and tables for each dataset. Lower values are better, with best in bold.
MIMIC-IV eICU
Column-wise $\downarrow$ Pair-wise $\downarrow$ High-order $\downarrow$ Column-wise $\downarrow$ Pair-wise $\downarrow$ High-order $\downarrow$
Model CDE I-CDE PCC I-PCC ER SMAPE CDE I-CDE PCC I-PCC ER SMAPE
Real - - - - 15.35 60.85 - - - - 44.05 43.42
SDV 0.11 0.54 0.26 0.26 49.32 103.85 0.13 0.61 0.19 0.19 76.34 101.97
RC-TGAN 0.26 0.54 0.18 0.28 38.21 97.26 0.34 0.59 0.18 0.21 69.08 111.25
ClavaDDPM 0.06 0.22 0.08 0.27 27.91 80.02 0.06 0.26 0.07 0.19 60.36 67.83
RawMed 0.04 0.05 0.04 0.10 19.69 57.31 0.05 0.08 0.06 0.10 45.58 48.42
Table 3: Clinical utility evaluation on MIMIC-IV and eICU datasets for downstream clinical tasks. Performance is measured by micro-averaged AUROC across 11 clinical prediction tasks (see Table 13 for individual task performance) under MEDS-TAB and GenHPF representations.
MIMIC-IV eICU
Model MEDS-TAB GenHPF MEDS-TAB GenHPF
Real $0.90\text{\scriptsize$\pm 0.06$}$ $0.82\text{\scriptsize$\pm 0.09$}$ $0.87\text{\scriptsize$\pm 0.08$}$ $0.80\text{\scriptsize$\pm 0.09$}$
SDV $0.46\text{\scriptsize$\pm 0.13$}$ $0.47\text{\scriptsize$\pm 0.13$}$ $0.46\text{\scriptsize$\pm 0.10$}$ $0.48\text{\scriptsize$\pm 0.08$}$
RC-TGAN $0.51\text{\scriptsize$\pm 0.14$}$ $0.51\text{\scriptsize$\pm 0.13$}$ $0.47\text{\scriptsize$\pm 0.13$}$ $0.48\text{\scriptsize$\pm 0.11$}$
ClavaDDPM $0.68\text{\scriptsize$\pm 0.19$}$ $0.64\text{\scriptsize$\pm 0.17$}$ $0.66\text{\scriptsize$\pm 0.14$}$ $0.63\text{\scriptsize$\pm 0.12$}$
RawMed $\mathbf{0.87}\text{\scriptsize$\pm 0.08$}$ $\mathbf{0.80}\text{\scriptsize$\pm 0.09$}$ $\mathbf{0.83}\text{\scriptsize$\pm 0.10$}$ $\mathbf{0.78}\text{\scriptsize$\pm 0.10$}$
Table 4: Temporal fidelity and privacy evaluation on MIMIC-IV and eICU datasets. Next Event Prediction F1 scores are averaged over three random seed runs with std. The best results are in bold.
Next Event Predict (F1) $\uparrow$ Time Gap $\downarrow$ Event Count $\downarrow$ MIA (Accuracy)
Model MIMIC-IV eICU MIMIC-IV eICU MIMIC-IV eICU MIMIC-IV eICU
Real $0.18\text{\scriptsize$\pm 0.06$}$ $0.30\text{\scriptsize$\pm 0.00$}$
SDV $0.05\text{\scriptsize$\pm 0.01$}$ $0.13\text{\scriptsize$\pm 0.00$}$ 0.76 0.50 0.46 0.36 0.499 0.502
RC-TGAN $0.02\text{\scriptsize$\pm 0.00$}$ $0.07\text{\scriptsize$\pm 0.01$}$ 0.43 0.39 0.04 0.03 0.500 0.500
ClavaDDPM $0.06\text{\scriptsize$\pm 0.00$}$ $0.12\text{\scriptsize$\pm 0.00$}$ 0.48 0.41 0.11 0.05 0.500 0.499
RawMed $\mathbf{0.16}\text{\scriptsize$\pm 0.00$}$ $\mathbf{0.25}\text{\scriptsize$\pm 0.03$}$ 0.01 0.03 0.02 0.05 0.498 0.497
Table 5: Results on MIMIC-IV and eICU datasets with a 6-hour observation window, comparing RealTabFormer (RTF) and RawMed. Lower is better, best in bold.
MIMIC-IV eICU
Metric RTF RawMed RTF RawMed
CDE $\downarrow$ 0.11 0.04 0.21 0.06
PCC $\downarrow$ 0.08 0.03 0.19 0.09
Time Gap $\downarrow$ 0.17 0.05 0.13 0.01
# Events $\downarrow$ 0.20 0.07 0.07 0.08
Table 6: Ablation study on MIMIC-IV dataset. Variants exclude RQ (Residual Quantization), Time Tok. (Time Tokenization), and Time Sep. (Time Separation). Lower values are better, best in bold.
Variants CDE PCC Time Gap # Events
Full Method 0.04 0.04 0.01 0.02
w/o RQ 0.05 0.04 0.05 0.13
w/o Time Tok. 0.04 0.04 0.04 0.12
w/o Time Sep. 0.07 0.07 0.51 0.40
5 Results
In single-table evaluations, as presented in Table 2, RawMed demonstrates superior performance across CDE and PCC metrics. Notably, ClavaDDPM exhibits performance closer to RawMed compared to SDV and RC-TGAN. However, in item-specific metrics, I-CDE and I-PCC, RawMed significantly outperforms ClavaDDPM. This underscores RawMed’s capability to accurately preserve the characteristics of individual clinical items. Additionally, in Predictive Similarity, RawMed achieves the lowest SMAPE and ER among baselines, closely matching real data, demonstrating semantic consistency with the original dataset alongside robust I-CDE and I-PCC performance. For a detailed results on table-wise and column-wise fidelity, refer to Table 11 and 12.
In multi-table time-series evaluations, the performance disparity between RawMed and baseline methods becomes more pronounced. For Clinical Utility, Table 3 shows RawMed’s AUROC scores closely match real data in downstream prediction tasks, surpassing ClavaDDPM, SDV, and RC-TGAN in both representations. Regarding temporal metrics, as shown in Table 4, RawMed excels in replicating inter-event interval distributions (Time Gap), achieving outstanding scores of 0.01–0.03 against baselines’ 0.41–0.76. For Event Count, RawMed records competitive scores of 0.02 (MIMIC-IV) and 0.05 (eICU), though RC-TGAN outperforms it in eICU (0.03). RawMed also excels in Next Event Prediction, capturing event sequence patterns more effectively than other methods. Additionally, RawMed exhibits minimal vulnerability to membership inference attacks (MIA), performing near random guessing levels, thus ensuring robust privacy preservation.
Collectively, these results highlight RawMed’s efficacy in generating multi-table time-series EHRs data that closely mirrors real data in terms of temporal precision, clinical utility, and privacy preservation. In contrast, baselines such as SDV, RC-TGAN, and ClavaDDPM underperform due to their lack of explicit temporal modeling and limited capability to model complex, large-scale EHR data.
Compressed vs. Non-compressed approaches
We compared RealTabFormer (RTF, non-compressed) and RawMed (compressed) to evaluate the effect of sequence compression on synthetic data quality.444 Experiments typically used a 12-hour observation window for training and synthesis, but both RTF and RawMed used a 6-hour window in this experiment due to RTF’s context-size limits. We synthesized about 20k samples to match RTF-generated sample counts, as generating the original dataset size was infeasible for RTF. RTF models data in the text space, handling maximum sequence lengths of 11.2k for MIMIC-IV and 3.1k for eICU. In contrast, RawMed operates in the latent space, reducing these to 1.8k (84% reduction) and 0.8k (74% reduction), respectively. As reported in Table 6, RawMed outperformed RTF in most metrics, including time gap metrics of 0.05 (MIMIC-IV) and 0.01 (eICU) compared to RTF’s 0.17 and 0.13, respectively. Thus, RawMed demonstrates superior data fidelity and temporal accuracy while significantly reducing sequence length.
(a) Real
(b) VQ Recon
[…]
(d) VQ (2x Code)
(e) RQ Recon
(f) RQ Synthetic
Figure 2: Comparison of VQ-VAE and RQ-VAE for the patientweight column in MIMIC-IV. Subfigures: (a) real data, (b) VQ reconstruction, (c) VQ with less compression, (d) VQ with doubled codebook, (e) RQ reconstruction, (f) RQ synthetic data. Shared x-axes with proportional y-axes.
6 Ablation Studies
RQ vs. VQ: Why RQ is more suitable
Although Vector Quantization (VQ) offers robust compression within the RawMed framework, it does not consistently achieve high-fidelity representation across all data types. In text-based EHR data, columns with low correlation and independent distributions (e.g., patientweight) are challenging to encode with VQ’s limited codebook, resulting in significant distortion of the original distribution upon reconstruction, as shown in Figure 2(b). Halving the compression ratio (Figure 2(c)) or doubling the codebook size (Figure 2(d))—note that these data are reconstructed, not generated—fails to fully resolve distortion in these numerical columns. In contrast, RawMed employs Residual Quantization (RQ) for multi-stage quantization, effectively preserving independent column distributions (Figure 2(e)–(f)). Table 14 shows that for patientweight in MIMIC-IV, VQ’s KS statistic (0.28) far exceeds RQ’s (0.09), confirming VQ’s greater distortion.
Ablation study of RawMed components
Table 6 evaluates RawMed variants with components removed, with lower metric values indicating better performance. The full method achieves optimal results across all metrics. Removing RQ (i.e., using VQ) degrades performance, increasing CDE to 0.05, Time Gap to 0.05, and Event Count to 0.13. Excluding Time Tokenization, which decomposes timestamps into digits with a 0–9 vocabulary, impairs time pattern recognition, raising Time Gap to 0.04 and Event Count to 0.12. Omitting Time Separation, which isolates timestamps for explicit temporal modeling, causes the largest performance drop, elevating CDE to 0.07, Time Gap to 0.51, and Event Count to 0.40. These results highlight the critical role of each component in RawMed’s temporal modeling efficacy.
Scalability
In this study, We conducted experiments primarily with a 12-hour observation window, also testing 6-hour and 24-hour windows to evaluate RawMed’s scalability. As shown in Table 15, most metrics remained stable across windows, though the KS statistics for the event count increased slightly for the 24-hour window. This demonstrates RawMed’s potential to generalize across different time scales.
7 Discussion
This study introduces RawMed, a framework for synthesizing multi-table time-series EHR data while retaining all columns and original values. By enabling hospitals to generate high-fidelity synthetic EHRs with minimal preprocessing and providing users with realistic data for flexible downstream tasks, RawMed can accelerate medical AI research while safeguarding patient privacy. However, several limitations remain with respect to fully encompassing the complexity of real-world medical settings. First, this work focuses on three primary tables, which leaves the question whether it can handle a significantly higher number (e.g., dozens) of tables. In addition, extended time-series data might require advanced compression or sampling methods. While the framework could potentially incorporate conditional generation (including static attributes such as gender or birth year), this study, as an initial approach, focuses on unconditional generation, leaving these specialized applications for future research. In subsequent work, we plan to expand the range of tables, implement conditional generation, and more fully integrate static features into the model design, thereby providing more realistic synthetic EHR data for a broader array of medical AI tasks.
Appendix A Feature definitions in EHR data
A.1 Feature definition
In electronic health record (EHR) data, tables (e.g., lab, medication, infusion) contain columns, some of which are item columns (e.g., itemid, drug, labname) that identify clinical events, such as which lab tests or medication administration. A feature is a processed representation of data derived from one or more columns, typically created by selecting, combining, or transforming column data, as described in the Introduction.11footnotemark: 1
To define a feature from EHR data, the following steps are typically applied:
  • Select a specific item: A specific item is selected to represent a clinical event. For example, in a laboratory table, selecting itemid = 12345 might isolate all instances of a “blood glucose” test.
  • Incorporate relevant categorical columns: Related categorical columns (e.g., uom, dose_unit, or route) further refine the feature definition. For instance, two rows with the same itemid but different uom values, such as “mg/dL” vs. “mmol/L”, would represent distinct features.
  • Include numerical columns: Numerical columns (e.g., valuenum, amount, or rate) provide the actual measurement or dosage values for the item. Typically, a feature can be defined by combining an item, a categorical column (e.g., uom), and a numerical column’s values. Each numerical column may also represent a distinct feature when it captures a unique aspect of a clinical event, such as a dosage rate versus a total amount.
For example, “blood glucose level” is defined by valuenum where itemid = 12345 and uom = “mg/dL”. While feature engineering is commonly used, this section focuses on the core structure of features, excluding feature engineering techniques (e.g., aggregation or transformation).
A.2 Feature count calculation
To calculate the number of features in EHR data, we define two approaches: the theoretical maximum (Possible Feature Count) and the practical estimate (Actual Feature Count). The data consists of a set of tables $\mathcal{E}$, such as medication, laboratory, or infusion tables. Each table $\epsilon\in\mathcal{E}$ includes categorical columns $C_{\epsilon,\text{cat}}$ and numerical columns $C_{\epsilon,\text{num}}$.
The following definitions are essential for calculating feature counts:
  • Distinct Items ($I_{\epsilon}$): The set of unique items in table $\epsilon$, representing distinct clinical event identifiers in an item column. For example, in a medication table, each item $i\in I_{\epsilon}$ corresponds to a drug name (e.g., paracetamol, aspirin).
  • Number of Unique Values in a Categorical Column ($|V_{i,c}|$): The count of unique values in categorical column $c\in C_{\epsilon,\text{cat}}$ (e.g., administration route) for item $i$. For instance, if paracetamol has administration routes “intravenous” and “oral,” then $|V_{i,\text{route}}|=2$.
  • Number of Unique Categorical Value Combinations ($|C_{\epsilon,\text{cat}}^{\text{unique}}(i)|$): The number of observed combinations of all categorical columns values for item $i$. For example, if paracetamol has three combinations of route and dose unit, then $|C_{\epsilon,\text{cat}}^{\text{unique}}(i)|=3$.
  • Number of Numerical Columns ($|C_{\epsilon,\text{num}}|$): The count of numerical columns (e.g., amount, rate) associated with each item-category combination.
Possible Feature Count
The Possible Feature Count assumes all possible combinations of categorical values exist for each item, providing a theoretical upper bound. It is calculated as:
Possible Feature $\text{Possible Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg{(}\sum_{i\in I_{\epsilon}}\Big{(}\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|\Big{)}\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg{)},$
where $\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|$ represents the number of possible categorical combinations, and $|C_{\epsilon,\text{num}}|+1$ accounts for the numerical columns plus a indicator feature representing the presence or absence of the categorical combination.
Actual Feature Count
The Actual Feature Count considers only the categorical combinations observed in the data, reflecting dependencies or sparsity that reduce the number of combinations. It is calculated as:
Actual Feature $\text{Actual Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg{(}\sum_{i\in I_{\epsilon}}\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg{)},$
where $\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|$ is the number of actual combinations, and $|C_{\epsilon,\text{num}}|+1$ accounts for numerical columns and the presence/absence feature.
For example, consider a medication table with 100 unique medications (drug), two categorical columns (dose_unit with 5 unique values and route with 3 unique values), and two numerical columns (amount and rate). If all $5\times 3=15$ combinations of dose_unit and route are valid, the features per medication are calculated as $15\times(2+1)=45$, resulting in a total of $100\times 45=4,500$ features. However, if only 10 combinations of dose_unit and route are present, the features per medication decrease to $10\times(2+1)=30$, and the total feature count reduces to $100\times 30=3,000$.
In summary, the Possible Feature Count provides a theoretical maximum, while the Actual Feature Count reflects the data’s sparsity and structure for a realistic estimate. This study used the Actual Feature Count to calculate the number of generated features (Table 1).
Appendix B Dataset statistics
B.1 Dataset selection criteria
To synthesize time-series electronic health records (EHR), we selected ICU patient datasets, specifically MIMIC-IV and eICU, for their high temporal resolution and continuous monitoring. MIMIC-IV, collected from Beth Israel Deaconess Medical Center, includes 94,458 ICU stays across 364,627 patients, while eICU, a multi-center dataset, comprises 200,859 ICU stays from 139,367 patients across U.S. hospitals. These datasets provide large cohorts and diverse clinical variables, making them established benchmarks for validating most prior EHR synthesis studies. We focused on time-series tables, labevents, prescriptions, and inputevents from MIMIC-IV, and lab, medication, and infusiondrug from eICU. These tables capture critical temporal changes in physiological states, treatment interventions, and drug administration, enabling the synthesis of complex clinical patterns. In contrast, emergency department (ED) and outpatient datasets, with irregular intervals and lower resolution, are typically less suitable for synthesizing the continuous temporal dynamics of patient trajectory.
However, to evaluate the model’s generalizability, we additionally utilized the MIMIC-IV-ED dataset to validate its ability to synthesize data in a distinct clinical setting. From MIMIC-IV-ED, we synthesized the vitalsign (time-series data such as heart rate and blood pressure), medrecon (static data on medications at admission), and pyxis (event-based data on medication dispensing via the Pyxis system) tables. By simultaneously generating time-series (vitalsign, pyxis) and non-time-series (medrecon) data, the model effectively synthesized complex patterns in ED data, demonstrating robustness across diverse clinical scenarios.
B.2 Data statistics
This study utilized the publicly available MIMIC-IV and eICU datasets, focusing on patients aged 18 years or older with ICU stays exceeding 12 hours. Key metrics, including patient counts, event frequencies, and row counts for the tables, are summarized in Table 7 Additionally, ablation studies exploring 6- and 24-hour observation windows for MIMIC-IV and eICU, and a 6-hour window for the MIMIC-IV-ED dataset in ED settings, are detailed in Table 8.
Table 7: Summary statistics for preprocessed MIMIC-IV and eICU datasets with a 12-hour observation window, including patient counts, event frequencies, and row and column counts for lab (labevents or lab), medication (prescriptions or medication), and input (inputevents or infusiondrug) data. Note: M denotes million.
Metric MIMIC-IV eICU
Number of Patients 62,704 143,812
Max. Events per Patient 243 114
Avg. Events per Patient 108.8 47.2
Total Rows 6.8M 6.8M
Rows: Lab 3.8M 4.6M
Rows: Medication 1.7M 1.6M
Rows: Input 1.4M 0.6M
Columns: Lab 9 6
Columns: Medication 9 6
Columns: Input 15 6
Table 8: Summary statistics for preprocessed MIMIC-IV-ED (6-hour observation window), MIMIC-IV, and eICU (6- and 24-hour observation windows) datasets. Note: M denotes million.
Metric MIMIC-IV-ED MIMIC-IV (6h) MIMIC-IV (24h) eICU (6h) eICU (24h)
Observation Window (hrs) 6 6 24 6 24
Number of Patients 144,790 63,903 51,357 132,202 119,641
Max. Events per Patient 34 165 366 79 179
Avg. Events per Patient 16.5 71.6 170.2 32.0 74.1
Total Rows 2.4M 4.6M 8.7M 4.2M 8.9M
B.3 Criteria for excluded columns
In this study, we aimed to utilize as many columns as possible but excluded specific columns for the following reasons and generated all others. First, we removed patient or prescription identifiers and unique hospital-specific codes (e.g., subject_id, orderid, pharmacy_id) that were not directly relevant to the clinical information used by the model.
Second, we excluded columns that contained multiple timestamps or offset columns conveying similar time information (e.g., charttime, starttime). While each timestamp could potentially have clinical value, we consolidated them into a single representative time field to reduce analytical complexity.
Finally, we excluded columns that potentially contained future information beyond the event time to avoid overestimating model performance by training on data not available at the actual prediction point. All remaining columns were included in the modeling process. Table 9 summarizes all excluded columns and indicates the primary reason for each exclusion.
Table 9: Columns excluded from each dataset table, categorized by the reason for exclusion: identifiers or codes, time-related columns, or columns containing potential future information.
Dataset Table Excluded Columns
Reason: Identifiers or Codes
MIMIC-IV labevents labevent_id, subject_id, specimen_id, order_provider_id
MIMIC-IV inputevents subject_id, caregiver_id
MIMIC-IV prescriptions subject_id, pharmacy_id, poe_id, poe_seq, order_provider_id, orderid, linkorderid
eICU lab labid
eICU infusiondrug infusiondrugid
eICU medication medicationid, drughiclseqno, gtc
Reason: Time Columns
MIMIC-IV labevents storetime
MIMIC-IV inputevents endtime, storetime
MIMIC-IV prescriptions stoptime
eICU lab labresultrevisedoffset
eICU medication drugorderoffset, drugstopoffset
Reason: Future Information Leakage
MIMIC-IV prescriptions statusdescription
eICU medication drugordercancelled
Appendix C Evaluation framework details
To rigorously evaluate the quality of synthetic multi-EHR data generated by RawMed, we define a comprehensive set of metrics that assess both single-table fidelity and multi-table temporal dynamics. Each metric is carefully selected to capture specific aspects of data quality, such as distributional similarity, inter-column dependencies, temporal fidelity, clinical utility, and privacy. Below, we provide detailed definitions, mathematical formulations, and justifications for each metric, supported by visualizations where applicable. Our evaluation pipeline is designed to be robust and tailored to the unique challenges of raw multi-table time-series EHR data, providing a thorough assessment of synthetic data quality.
C.1 Single-table evaluation
Single-table evaluation focuses on the fidelity of synthetic data within individual tables, treating each row as an independent instance. We combine low-order and high-order metrics to capture marginal distributions and complex inter-column dependencies. Below, we detail each metric, its formulation, and its relevance to EHR data evaluation.
  • Column-wise Density Estimation (CDE): CDE measures the similarity of distributions between real and synthetic data for each column. For numeric columns, we use the Kolmogorov-Smirnov (KS) statistic, defined as:
    $\text{KS}=\sup_{x}|F_{r}(x)-F_{s}(x)|,$
    where $F_{r}(x)$ and $F_{s}(x)$ are the cumulative distribution functions (CDFs) of the real and synthetic data, respectively. The KS statistic ranges from $[0,1]$, with lower values indicating higher similarity. For categorical columns, we use the Jensen-Shannon (JS) divergence, defined as:
    $\text{JS}(P_{r}\|P_{s})=\frac{1}{2}\text{KL}(P_{r}\|M)+\frac{1}{2}\text{KL}(P_{s}\|M),$
    where $P_{r}$ and $P_{s}$ are the probability distributions of the real and synthetic data, $M=\frac{1}{2}(P_{r}+P_{s})$, and KL is the Kullback-Leibler divergence. JS ranges from $[0,1]$, with lower values indicating greater similarity.
    • Justification: CDE is a fundamental metric for comparing distributional similarity across tabular data. It effectively captures the distributional properties of numerical columns (e.g., laboratory values) and categorical columns (e.g., prescribed drug names) in EHR data. This ensures robust evaluation across diverse column types, critical for synthetic data validity.
  • Item-specific Column-wise Density Estimation (I-CDE): I-CDE extends CDE by evaluating distributional fidelity for specific clinical items (e.g., creatinine, glucose) within columns that aggregate multiple item types (e.g., a value column linked to an itemid). For each item, we filter the data and apply the same KS or JS metrics as in CDE. To ensure robustness, we exclude items with fewer than 0.1% of total records (e.g., lab tests occurring 1–2 times in 3.8 million events) to avoid skewed results due to extremely rare events.
    • Justification: I-CDE addresses the heterogeneity of EHR data, where a single column may represent diverse clinical entities (e.g., 2,000 drugs). By ensuring synthetic data preserves item-specific distributions, I-CDE enhances clinical validity.
  • Pairwise Column Correlation (PCC): PCC assesses the fidelity of inter-column dependencies by comparing correlation matrices of real and synthetic data. For numeric-numeric pairs, we use Pearson’s correlation coefficient ($\rho\in[-1,1]$). For categorical-categorical pairs, we use Theil’s U statistic ($U\in[0,1]$), which measures conditional entropy. For categorical-numeric pairs, we use the correlation ratio ($\eta\in[0,1]$), which quantifies the variance explained by the categorical variable. The mean absolute difference ($\mu_{\mathrm{abs}}$) between real and synthetic correlation matrices is computed as:
    $\mu_{\mathrm{abs}}=\frac{1}{N}\sum_{i,j}|\text{Corr}_{r}(i,j)-\text{Corr}_{s}(i,j)|,$
    where $\text{Corr}_{r}$ and $\text{Corr}_{s}$ are the correlation matrices for real and synthetic data, and $N$ is the number of matrix elements.
    • Justification: PCC captures dependencies between clinical variables (e.g., heart rate and blood pressure), essential for realistic synthetic EHR data. The use of multiple correlation measures ensures applicability to mixed data types in EHRs.
  • Item-specific Pairwise Column Correlation (I-PCC): I-PCC extends PCC by computing correlation matrices for each item-specific subset of the data (e.g., all records for a specific itemid). The $\mu_{\mathrm{abs}}$ is calculated for each item’s matrix and averaged across items. As with I-CDE, items with fewer than 0.1% of records are excluded.
    • Justification: I-PCC evaluates item-specific inter-column dependencies, ensuring that synthetic data captures relationships within subsets of EHR data (e.g., correlations between dosage and frequency for a specific drug). This is critical for maintaining clinical relevance.
  • Predictive Similarity: Predictive Similarity evaluates the ability of synthetic data to capture high-order, non-linear dependencies by training an XGBoost model to predict each column using the remaining columns as features. Models are trained on synthetic data and evaluated on real test data. For numeric targets, we use Symmetric Mean Absolute Percentage Error (SMAPE), defined as:
    $\text{SMAPE}=\frac{100}{n}\sum_{i=1}^{n}\frac{|y_{i}-\hat{y}_{i}|}{(|y_{i}|+|\hat{y}_{i}|)/2},$
    where $y_{i}$ and $\hat{y}_{i}$ are the real and predicted values, and SMAPE ranges from [0,200]. For categorical targets, we use the classification error rate (ER), defined as:
    $\text{ER}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(y_{i}\neq\hat{y}_{i})\times 100,$
    where ER ranges from [0,100]. A smaller performance gap between models trained on synthetic versus real data indicates better capture of dependencies.
    • Justification: Predictive similarity captures complex relationships that low-order metrics may not detect, indirectly assessing semantic consistency. While Mean Squared Error (MSE) and Mean Absolute Error (MAE) are commonly used regression metrics, their scale-dependency led to the adoption of SMAPE, which is bounded between 0 and 200 for consistent evaluation across varying ranges.
C.2 Time-series multi-table evaluation
Multi-table EHR data involve time-series events across multiple tables, linked by a primary key (e.g., stay_id). Our evaluation metrics preserve this structure and assess temporal fidelity, clinical utility, and privacy. Below, we detail each metric and its rationale.
  • Clinical Utility: To assess clinical utility, we evaluate the performance of synthetic data on downstream predictive tasks. We defined 11 clinical prediction tasks (e.g., predicting creatinine and hemoglobin levels), following the task definitions from genhpf . Details are provided in Table 10. To address variability in the number and types of events across patients, we employ two representation methods, GenHPF and MEDS-TAB, as detailed in E.2.1. For each task, the synthetic data is split into two segments over an observation window of length $T$ (e.g., $T=12$ hours. The first half (e.g., the first 6 hours) serves as input for the predictive model, with predictions made at the midpoint ($T/2$). For multi-class tasks, the label is the last value of the lab measurement in the second half (e.g., the subsequent 6 hours). For binary tasks, the label indicates whether a medication event occurs in the second half. Models are trained on synthetic data and tested on real data, with performance measured using the Area Under the Receiver Operating Characteristic Curve (AUROC). Higher AUROC scores indicate better utility for clinical applications.
    • Justification: Clinical utility reflects the practical value of synthetic data for real-world EHR applications. Using two representation methods ensures robustness across diverse modeling approaches.
  • Membership Inference Attacks (MIA): MIA assesses privacy leakage by attempting to infer whether a sample was part of the training dataset. We compute distances (e.g., Hamming distance) between synthetic samples and the real training set, using a threshold-based classifier to predict membership. The attack’s success rate (accuracy) indicates privacy risk, with lower values reflecting better privacy protection. This is conducted using the GenHPF representation.
    • Justification: MIA is a standard metric for evaluating privacy in synthetic EHR data, as it measures the risk of re-identifying individuals from the training set, a critical concern given EHRs’ sensitive patient information.
  • Time Gap: Time Gap evaluates the similarity of time interval distributions between consecutive patient events, using the KS statistic. In EHR data, simultaneous events (zero intervals) are common due to concurrent clinical recordings in clinical workflows or temporal resolution constraints of EHR systems. To address this, we compute KS statistics both including and excluding zero intervals, with results excluding zero intervals reported in Table 4. CDFs for both cases are visualized in Figure 4: panels (a)–(b) show distributions including simultaneous events, and panels (c)–(d) show distributions excluding them. Additionally, absolute event time distributions (from admission to event occurrence) are presented in Figure 4(e)–(f).
    • Justification: Time Gap ensures synthetic data accurately captures the timing of clinical events, essential for realistic patient trajectories in time-sensitive settings like ICUs. By evaluating both zero and non-zero intervals, it accounts for EHR-specific recording patterns, while the KS statistic provides a robust, non-parametric measure of temporal fidelity.
  • Event Count: Event Count evaluates the similarity in the distribution of the number of clinical events per patient between synthetic and real EHR data, employing the (KS statistic to compare these distributions. The CDFs of event counts for both synthetic and real data are visualized in Figure 4(g)–(h).
    • Justification: Event Count verifies that synthetic data replicates the frequency of clinical events per patient, a key characteristic of patient trajectories.
  • Next Event Prediction: Next Event Prediction evaluates the synthetic data’s ability to capture temporal sequence dynamics. An LSTM-based model (lstm, ) predicts the next event’s item or drug name, formulated as a multi-label classification task to handle concurrent events. The model is trained on synthetic data and evaluated on real test data using the F1 score, which balances precision and recall for multi-label predictions. Higher F1 scores indicate better sequence modeling.
    • Justification: This metric captures the sequential patterns of clinical events, essential for modeling patient trajectories. The LSTM’s ability to model long-term dependencies suits the complex temporal structure of EHRs.
C.3 Visualizations and results
To provide a comprehensive view of synthetic data quality, we include the following figures and tables:
  • Figure 3: Correlation matrix heatmap showing differences in Pairwise Column Correlation (PCC) between real and synthetic data.
  • Figure 4: Cumulative Distribution Function (CDF) plots for Time Gap (including and excluding zero intervals), absolute event time, and event count distributions.
  • Table 2: Overall single-table evaluation results, averaged across tables and columns for each dataset.
  • Tables 11, 12: Column-wise evaluation results, including CDE, I-CDE, SMAPE, ER gaps for predictive similarity, and null ratio of real and synthetic data.
  • Table 3: AUROC comparisons for clinical utility tasks across datasets.
  • Table 10: AUROC results for clinical utility tasks, detailing performance for each task.
  • Table 4: Accuracy for Membership Inference Attacks (MIA), F1 scores for Next Event Prediction, and KS statistics for Time Gap and Event Count.
By combining rigorous metric definitions, mathematical formulations, and extensive visualizations, our evaluation pipeline offers a thorough and principled approach to assessing synthetic multi-table EHR data. This comprehensive framework not only addresses existing gaps in evaluation methodologies but also sets a new standard for evaluating time-series data in healthcare applications.
Appendix D RawMed framework details
D.1 Architecture details
This section details the architectural components of the RawMed framework, consisting of two distinct modules: event compression and inter-event temporal modeling, providing specifics beyond the main text.
The event compression module transforms serialized clinical event text into compact latent representations, leveraging a Vector Quantized Variational AutoEncoder (VQ-VAE) vqvae or Residual Quantization (RQ) framework rqvae , both implemented using 1D convolutional neural networks (CNNs). The processing pipeline begins with the construction of input embeddings, followed by encoding, quantization, and decoding stages.
The input embedding, denoted $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, integrates three components adapted from an established text-based EHR prediction framework, GenHPF, to ensure robust representation of clinical data. The textual embedding ($\mathbf{x}_{\text{text},i}^{p}$) employs framework-derived embeddings for serialized event text, such as “lab item Glucose value 95 uom mg/dL”, tokenized using a Bio+Clinical BERT tokenizer bioclinicalbert . The type embedding ($\mathbf{x}_{\text{type},i}^{p}$) assigns categorical labels to tokens, distinguishing table names, column names, and values to capture the relational structure of the data. The digit-place embedding ($\mathbf{x}_{\text{dpe},i}^{p}$) enhances numeric tokens by tokenizing digits with spaces (e.g., “123.45” as “1 2 3 . 4 5”) and assigning position-specific types for integer and decimal places. These components are combined as $\mathbf{x}_{i}^{p}=\mathbf{x}_{\text{text},i}^{p}+\mathbf{x}_{\text{type},i}^{p}+\mathbf{x}_{\text{dpe},i}^{p}$. The main text adopts the notation $\mathbf{x}_{i}^{p}=\mathbf{x}_{\text{text},i}^{p}$ for simplicity, though the full embedding incorporates type and digit-place embeddings to enhance structural and numeric fidelity.
The encoder (Enc) processes the input embedding through five CNN layers, each with a kernel size of 5, stride of 2, and padding of 2, followed by batch normalization and ReLU activation. This configuration ensures a receptive field sufficient to encompass the entire event sequence of length $L=128$, capturing contextual dependencies. The encoder outputs a latent representation, $\hat{\mathbf{z}}_{i}^{p}\in\mathbb{R}^{L_{z}\times F_{z}}$, which is subsequently quantized.
The decoder (Dec) reconstructs the text, type, and digit-position embeddings ($\hat{\mathbf{x}}_{\text{text},i}^{p},\hat{\mathbf{x}}_{\text{type},i}^{p},\hat{\mathbf{x}}_{\text{dpe},i}^{p}$) using transposed CNN layers. The training objective minimizes reconstruction losses for all three embedding components, ensuring faithful recovery of text, type, and digit-position information.
The temporal modeling module, implemented as TempoTransformer, is a Transformer-based architecture for autoregressive prediction of quantized event sequences interleaved with tokenized timestamps. The model employs causal attention to ensure predictions depend only on prior tokens. Positional encodings provide sequence order information, and the model is trained from scratch, as the autoregressive target is the quantized codebook, obviating the need for fine-tuning from pretrained weights.
D.2 Postprocessing
This appendix elaborates on the postprocessing steps applied to convert synthetic patient trajectories into relational tables, aligning with the structural and semantic consistency of real data. The core steps of event-level verification and patient-level validation, which ensure valid event formats and temporal consistency, are implemented in Algorithm 1. Here, we detail the additional postprocessing steps that enforce column-specific constraints, as referenced in Section 3.4, to refine the synthetic tables’ statistical fidelity.
After validated sequences are converted into relational tables, further postprocessing ensures compliance with column-specific constraints, addressing potential errors from the generative process (e.g., quantization or autoregressive sampling). The following steps are applied:
  • Numeric Columns: Values are filtered to lie within the minimum and maximum ranges observed in the real data, applied at both the column and item levels. Non-compliant events (e.g., a glucose value outside the real data’s range) are removed.
  • Categorical Columns: Invalid values not present in the permissible values for each column are replaced with the closest valid value, determined using Levenshtein distance similarity. A distance threshold ensures that only sufficiently similar replacements are accepted; otherwise, the value is flagged as invalid.
  • Patient-Level Filtering: If any events contains a numeric value outside the valid range or a categorical value exceeding the Levenshtein distance threshold, the entire patient sample associated with that row is removed to maintain data integrity.
These steps ensure that the final relational tables adhere to the structural and statistical characteristics of the real data, removing outliers and ensuring both numerical and categorical validity.
Input: Text-based events $E$, timestamps $\mathcal{T}$, valid column names $\mathcal{C}$, valid table names $\mathcal{E}$, observation window $\mathcal{T}_{\text{max}}$
Output: Relational tables $\mathcal{R}$
Initialize $\mathcal{R}\leftarrow\emptyset$ (dictionary of tables)
for each event $e_{i}\in E$ with timestamp $t_{i}\in\mathcal{T}$ do
       Parse $e_{i}$ into table name $\tau$ and column-value pairs $\{(c_{i,j},v_{i,j})\}$
      if $\tau\notin\mathcal{E}$ then
             Discard $e_{i}$
            Continue
       end if
      if $e_{i}$ does not conform to column-value pair format then
             Discard $e_{i}$
            Continue
       end if
      for each $(c_{i,j},v_{i,j})$ in $\{(c_{i,j},v_{i,j})\}$ do
             if $c_{i,j}\notin\mathcal{C}$ then
                   $c_{i,j}\leftarrow\arg\min_{c\in\mathcal{C}}\text{Levenshtein}(c_{i,j},c)$
             end if
            if $c_{i,j}$ is numeric and $v_{i,j}$ contains non-numeric characters then
                   $v_{i,j}\leftarrow\text{RemoveNonNumeric}(v_{i,j})$
             end if
       end for
      Add $(t_{i},\{(c_{i,j},v_{i,j})\})$ to $\mathcal{R}[\tau]$
end for
for each table $\mathcal{R}[\tau]$ in $\mathcal{R}$ do
       Sort events by timestamp $t_{i}$ in ascending order
      for each event $(t_{i},\{(c_{i,j},v_{i,j})\})$ in $\mathcal{R}[\tau]$ do
             if $i>1$ and $t_{i}<t_{i-1}$ or $t_{i}>\mathcal{T}_{\text{max}}$ then
                   Discard $(t_{i},\{(c_{i,j},v_{i,j})\})$ and all events $(t_{k},\{(c_{k,j},v_{k,j})\})$ for $k\geq i$
                  Break
             end if
       end for
end for
return $\mathcal{R}$
Algorithm 1 Postprocessing for Relational Table Construction
Appendix E Training details and hyperparameters
E.1 Generative models
  • RawMed The framework employs a two-stage training process: the event compression module is trained first to generate quantized representations, followed by the inter-event temporal modeling module to predict sequences autoregressively.
    • Event-level compression: The event compression module uses the AdamW optimizer (learning rate: 5e-4, weight decay: 0.01), processing batches of 4096 events for up to 200 epochs, with early stopping after 10 epochs of stagnant validation accuracy. The loss function combines reconstruction losses for text, type, and digit-place embeddings with a commitment loss (commitment cost: 1.0). An EMA decay factor of 0.8 is applied for codebook updates in VQ-VAE training. A dropout rate of 0.2 is used for regularization. The input embedding has sequence length $L=128$ and embedding dimension $F=256$, compressed to a latent representation with $L_{z}=4$ and $F_{z}=256$. The codebook contains $K=1024$ entries, with residual quantization (RQ) using a depth of $D=2$, yielding a $4\times 2$ code representation. Training is performed on a single NVIDIA A6000 GPU, completing in under 24 hours.
    • Temporal modeling between events: The temporal modeling module uses the AdamW optimizer (learning rate: 3e-4, weight decay: 0.01), with batches of 32 sequences for up to 200 epochs, with early stopping after 10 epochs without improvement. The TempoTransformer consists of 12 layers, each with 8 attention heads, a hidden dimension of 512, and a feed-forward dimension of 2048. The input sequence includes time tokens ($\tau_{i}^{p}\in\{0,\dots,9\}^{2}$, length 2), event tokens ($k_{i}^{p}\in[1024]^{4\times 2}$, codebook size 1024), and special tokens (<SOS>, <EOS>, <PAD>), resulting in a vocabulary size of 1037. Training was conducted on three NVIDIA A6000 GPUs for MIMIC and on a single NVIDIA A6000 GPU for eICU, both completing in under 48 hours.
    • Generation: Data generation uses top-$k$ sampling with $k=250$ for MIMIC-IV and $k=150$ for eICU.
  • RealTabFormer realtabformer : Text-based autoregressive model for generating relational database. This model involves fine-tuning the Meta-LLaMA-3.1-8B model and a generation step using top-p sampling.
    • Fine-tuning: The Meta-LLaMA-3.1-8B model dubey2024llama was fine-tuned using QLoRA dettmers2024qlora (rank $r$=128, LoRA scaling factor $\alpha$=128, LoRA dropout=0.05) and Flash Attention-2 for memory optimization. Training used a maximum sequence length of 11,240 tokens, matching the input data. The Adafactor optimizer was applied with a learning rate of 2e-4, weight decay of 0.001, a batch size of 1, gradient accumulation (step size=4), and gradient checkpointing. Training ran for 2 epochs on a single RTX 3090 GPU, taking approximately 10 days for MIMIC-IV and 20 days for eICU.
    • Generation: Used the fine-tuned Meta-LLaMA-3.1-8B model with top-p sampling, setting the temperature to 1 and a threshold of 0.7.
  • SDV sdv : The Synthetic Data Vault (SDV) uses the Hierarchical ML Algorithm (HMA) Synthesizer to generate synthetic relational data, modeling individual tables with Gaussian Copulas while preserving parent-child relationships.The SDV framework’s default hyperparameters are applied for training.
  • RC-TGANrctgan A GAN-based model designed to generate synthetic relational databases by modeling parent-child relationships.
    • Training and generation: Uses default hyperparameters from the official repository, with batch size set to 500,000, the maximum feasible value for the NVIDIA RTX 3090 GPU (default: 500). Other hyperparameters include embedding dimension: 128, generator dimensions: (256, 256), discriminator dimensions: (256, 256), epochs: 300, discriminator steps: 1, generator learning rate: 2e-4, and discriminator learning rate: 2e-4. Hyperparameter search tested discriminator steps ([1, 3]), generator learning rate ([1e-4, 2e-4, 4e-4]), and discriminator learning rate ([1e-4, 2e-4, 4e-4]), but defaults performed best. Training took less than 10 hours on a single NVIDIA RTX 3090 GPU without early stopping.
  • ClavaDDPMclavaddpm A diffusion-based model for multi-table data generation, Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM) leverage Denoising Diffusion Probabilistic Models (DDPMs) to model complex tabular data distributions. By using clustering labels as intermediaries, ClavaDDPM captures long-range dependencies across interconnected tables, particularly through foreign key constraints.
    • Training and generation: We adopted the default hyperparameters from the official ClavaDDPM GitHub, featuring a diffusion model with six MLP layers ([512, 1024, 1024, 1024, 1024, 512]), a learning rate of 6e-4, and a batch size of 4096 for 100,000 iterations with a cosine scheduler. The model employs 2000 timesteps with MSE loss and uses 50 clusters for clustering. Training was completed in under 3 hours on a single NVIDIA A6000 GPU.
E.2 Predictive models for evaluation
We use the following predictive models to assess the quality of synthetic data:
E.2.1 Clinical utility
  • GenHPF genhpf : GenHPF addresses the heterogeneity of multi-table EHR data by transforming all patient events into a single hierarchical textual sequence. This method sequentially lists events such as medications, infusions, and lab results in chronological order, requiring minimal preprocessing. GenHPF enables predictive models to capture the full complexity of patient records, making it suitable for multi-task learning without extensive domain-specific feature engineering. Its flexibility supports applications across diverse EHR systems, enhancing model generalizability for downstream predictive tasks.
    • Training: GenHPF was trained with an embedding dimension of 128, 4 attention heads, and 2 transformer layers. It uses a batch size of 64, a dropout rate of 0.1, and a learning rate of 5e-5 for 50 epochs. Training occurred on a single NVIDIA A6000 GPU, targeting multi-task prediction for EHR lab tests and medications, with early stopping after 10 epochs of no improvement.
  • MEDS-TAB medstab : MEDS-TAB is an automated tabularization tool that converts Medical Event Data Standard (MEDS)-formatted EHR data into standardized tabular representations by aggregating events into fixed time intervals. It processes irregularly sampled time-series data, supporting various aggregation functions (e.g., sum, count, mean) and window sizes to handle diverse data types, including static codes, numerical values, and time-series events. With high scalability and efficiency—demonstrated by processing 500 patients from MIMIC-IV in 16 seconds with 1,410MB memory usage—MEDS-TAB simplifies predictive modeling by providing structured inputs compatible with machine learning algorithms like XGBoost, achieving competitive performance in clinical tasks.
    • Training: MEDS-TAB was trained using the default hyperparameters from the official MEDS-TAB GitHub repository, employing XGBoost for predictive modeling. In contrast to GenHPF, it trains on individual tasks rather than multi-task learning.
E.2.2 Predictive similarity
For predictive similarity evaluation, the XGBoost model was configured with 100 estimators, a subsample ratio of 0.9, and a maximum bin size of 256. Depending on the task, objectives were set to binary:logistic for binary classification, multi:softmax for multiclass classification, or reg:squarederror for regression, with evaluation metrics logloss, mlogloss, or rmse, respectively.
E.2.3 Next event prediction
For the Next Event Prediction task, we employ a single-layer LSTM with 128 hidden units, with input and output sizes equal to the number of classes. A sigmoid activation function is used for multi-label classification. The model is trained with the Adam optimizer at a learning rate of 0.001, employing a weighted binary cross-entropy loss (with inverse log-frequency weights normalized to a mean of 1), a batch size of 256, and 50 epochs. A classification threshold of 0.1 is applied to enhance recall for imbalanced classes.
Table 10: Summary of 11 ICU prediction tasks using synthetic EHR data. Multi-class tasks predict binned lab values, while binary tasks classify whether specific medications are prescribed. Predictions are made at the midpoint ($T/2$) of an observation window of length $T$ (e.g., $T=12$ hours) to forecast the last lab value or medication event occurrence in the second half.
Task Name Source Task Type
Creatinine (Cr) Lab Multi-class
Platelets (Plt) Lab Multi-class
White Blood Cells (WBC) Lab Multi-class
Hemoglobin (Hb) Lab Multi-class
Bicarbonate (HCO3) Lab Multi-class
Sodium (Na) Lab Multi-class
Magnesium Sulfate (MgSO4) Medication Binary
Heparin (Hep) Medication Binary
Potassium Chloride (KCl) Medication Binary
Norepinephrine (NE) Input Binary
Propofol (Prop) Input Binary
Appendix F Additional experimental results
F.1 Column-wise fidelity metrics
We report column-wise fidelity metrics for RawMed-generated data on the MIMIC-IV and eICU datasets, evaluating distributional similarity and predictive performance across categorical and numerical columns. Notably, RawMed-generated null ratios closely align with real data, as detailed in Tables 11 and 12.
Table 11: Column-wise fidelity for MIMIC-IV. Columns are prefixed with “L” (laboratory measurements), “I” (infusions), or “M” (medications). Types “C” and “N” denote categorical and numerical columns, respectively. TRTR (Train-on-Real-Test-on-Real) and TSTR (Train-on-Synthetic-Test-on-Real) denote error rates for categorical columns and SMAPE for numerical columns, reported as mean ± standard deviation (std) across three random seeds, assessing predictive similarity task performance.
Col. Name Type CDE $\downarrow$ I-CDE $\downarrow$ Real Null (%) Syn Null (%) TRTR TSTR
L/itemid C 0.06 - 0.00 0.00 66.08 ±$\pm 24.72$ 78.87 ±$\pm 23.05$
L/value C 0.07 0.14 5.78 4.51 0.00 ±$\pm 0.00$ 0.77 ±$\pm 0.02$
L/valueuom C 0.04 0.03 13.78 12.49 0.07 ±$\pm 0.00$ 2.50 ±$\pm 0.02$
L/flag C 0.00 0.00 62.27 64.18 - -
L/priority C 0.03 0.04 27.34 29.48 44.16 ±$\pm 0.08$ 44.32 ±$\pm 0.19$
L/comments C 0.11 0.22 83.41 86.15 57.34 ±$\pm 0.61$ 63.52 ±$\pm 2.04$
L/valuenum N 0.01 0.06 11.13 9.46 116.64 ±$\pm 0.76$ 86.37 ±$\pm 1.99$
L/ref_range_lower N 0.02 0.01 19.04 17.15 60.86 ±$\pm 0.21$ 63.10 ±$\pm 0.20$
L/ref_range_upper N 0.01 0.01 19.04 17.15 30.45 ±$\pm 0.66$ 30.01 ±$\pm 0.36$
I/itemid C 0.09 - 0.00 0.00 28.17 ±$\pm 3.03$ 29.19 ±$\pm 0.13$
I/amountuom C 0.02 0.02 0.00 0.00 0.07 ±$\pm 0.00$ 0.13 ±$\pm 0.00$
I/rateuom C 0.05 0.05 37.49 39.54 0.02 ±$\pm 0.00$ 0.07 ±$\pm 0.00$
I/ordercategoryname C 0.03 0.03 0.00 0.00 0.16 ±$\pm 0.00$ 0.19 ±$\pm 0.00$
I/secondaryordercategoryname C 0.02 0.01 25.82 26.57 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
I/ordercomponenttypedescription C 0.01 0.01 0.00 0.00 0.00 ±$\pm 0.00$ 0.01 ±$\pm 0.00$
I/ordercategorydescription C 0.02 0.03 0.00 0.00 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
I/totalamountuom C 0.00 0.00 12.35 13.59 - -
I/isopenbag C 0.01 0.01 0.00 0.00 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
I/amount N 0.03 0.08 0.00 0.05 88.26 ±$\pm 0.87$ 87.75 ±$\pm 2.87$
I/rate N 0.04 0.13 37.49 39.54 36.47 ±$\pm 1.69$ 62.34 ±$\pm 1.11$
I/patientweight N 0.09 - 0.00 0.00 21.23 ±$\pm 0.01$ 21.78 ±$\pm 0.02$
I/totalamount N 0.02 0.07 12.38 13.63 6.77 ±$\pm 0.08$ 8.81 ±$\pm 0.02$
I/originalamount N 0.03 0.12 0.17 0.10 28.00 ±$\pm 1.23$ 30.78 ±$\pm 0.34$
I/originalrate N 0.03 0.09 5.83 4.46 41.72 ±$\pm 3.88$ 48.13 ±$\pm 2.02$
M/drug C 0.08 - 0.00 0.00 31.00 ±$\pm 0.38$ 40.53 ±$\pm 1.01$
M/drug_type C 0.00 0.00 0.00 0.00 0.82 ±$\pm 0.07$ 1.41 ±$\pm 0.04$
M/prod_strength C 0.08 0.05 0.09 0.04 17.69 ±$\pm 0.63$ 37.79 ±$\pm 0.43$
M/dose_unit_rx C 0.03 0.02 0.09 0.04 5.93 ±$\pm 0.11$ 13.06 ±$\pm 0.28$
M/form_unit_disp C 0.03 0.04 0.09 0.05 8.16 ±$\pm 0.21$ 21.07 ±$\pm 0.45$
M/route C 0.04 0.04 0.09 0.04 31.99 ±$\pm 0.49$ 40.59 ±$\pm 0.30$
M/doses_per_24_hrs N 0.02 0.06 59.97 61.97 60.30 ±$\pm 0.11$ 60.36 ±$\pm 0.12$
M/dose_val_rx N 0.02 0.04 6.13 6.38 109.53 ±$\pm 0.20$ 105.76 ±$\pm 0.57$
M/form_val_disp N 0.02 0.03 5.59 5.99 129.93 ±$\pm 13.27$ 82.49 ±$\pm 4.43$
Table 12: Column-wise fidelity for eICU. Columns are prefixed with “L” (laboratory measurements), “I” (infusions), or “M” (medications). Types “C” and “N” denote categorical and numerical columns, respectively. TRTR (Train-on-Real-Test-on-Real) and TSTR (Train-on-Synthetic-Test-on-Real) denote error rates for categorical columns and SMAPE for numerical columns, reported as mean ± standard deviation (std) across three random seeds, assessing predictive similarity task performance.
Col. Name Type CDE $\downarrow$ I-CDE $\downarrow$ Real Null (%) Syn Null (%) TRTR TSTR
L/labname C 0.04 - 0.00 0.00 52.57 ±$\pm 3.92$ 53.10 ±$\pm 12.26$
L/labmeasurenamesystem C 0.03 0.07 5.24 4.82 0.07 ±$\pm 0.02$ 0.59 ±$\pm 0.07$
L/labmeasurenameinterface C 0.04 0.14 6.37 5.59 54.97 ±$\pm 6.30$ 49.94 ±$\pm 0.49$
L/labtypeid C 0.03 0.06 0.00 0.00 0.00 ±$\pm 0.00$ 0.28 ±$\pm 0.04$
L/labresult N 0.01 0.11 0.93 0.64 14.68 ±$\pm 0.55$ 24.54 ±$\pm 0.65$
L/labresulttext N 0.01 0.12 0.75 0.54 19.72 ±$\pm 0.46$ 23.40 ±$\pm 0.35$
I/drugname C 0.16 - 0.00 0.00 57.38 ±$\pm 0.32$ 60.48 ±$\pm 0.20$
I/infusionrate N 0.03 0.43 55.54 63.14 77.70 ±$\pm 0.27$ 77.95 ±$\pm 1.08$
I/drugamount N 0.06 0.32 67.48 74.43 71.46 ±$\pm 1.57$ 69.29 ±$\pm 0.18$
I/volumeoffluid N 0.03 0.23 67.47 74.37 14.80 ±$\pm 0.47$ 18.11 ±$\pm 0.18$
I/patientweight N 0.03 - 90.91 95.91 12.49 ±$\pm 0.20$ 33.41 ±$\pm 0.44$
I/drugrate N 0.03 0.36 0.46 0.55 93.12 ±$\pm 0.66$ 92.25 ±$\pm 0.25$
M/drugname C 0.10 - 32.53 29.65 73.75 ±$\pm 0.17$ 81.10 ±$\pm 0.47$
M/drugivadmixture C 0.02 0.03 0.00 0.00 14.49 ±$\pm 0.07$ 16.36 ±$\pm 0.01$
M/dosage C 0.07 0.15 11.13 11.38 80.19 ±$\pm 0.08$ 81.74 ±$\pm 0.17$
M/routeadmin C 0.07 0.11 0.00 0.01 55.23 ±$\pm 0.15$ 57.59 ±$\pm 0.09$
M/frequency C 0.08 0.22 12.60 11.31 85.52 ±$\pm 0.19$ 88.22 ±$\pm 0.10$
M/prn C 0.00 0.06 0.01 0.08 10.40 ±$\pm 0.12$ 11.97 ±$\pm 0.10$
F.2 Task-specific predictive performance
We report task-wise predictive performance for clinical tasks using RawMed-generated data on the MIMIC-IV and eICU datasets, comparing Train-on-Real-Test-on-Real (TRTR) and Train-on-Synthetic-Test-on-Real (TSTR) metrics, as detailed in Table 13.
Table 13: Clinical task utility evaluation. Train-on-Synthetic-Test-on-Real (TSTR) performance for MIMIC-IV and eICU datasets, reported as mean ± standard deviation (std) across three random seeds. TRTR denotes Train-on-Real-Test-on-Real, and TSTR denotes Train-on-Synthetic-Test-on-Real. GenHPF and MEDS-TAB are synthetic data generation methods. Tasks include laboratory value predictions (Creatinine: Cr, Platelets: Plt, White Blood Cells: WBC, Hemoglobin: Hb, Bicarbonate: HCO3, Sodium: Na) and medication/input predictions (Magnesium Sulfate: MgSO4, Heparin: Hep, Potassium Chloride: KCl, Norepinephrine: NE, Propofol: Prop).
MIMIC-IV eICU
GenHPF MEDS-TAB GenHPF MEDS-TAB
Task TRTR TSTR TRTR TSTR TRTR TSTR TRTR TSTR
Cr 0.89 ±$\pm 0.00$ 0.88 ±$\pm 0.00$ 0.97 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.86 ±$\pm 0.00$ 0.83 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.89 ±$\pm 0.00$
Plt 0.88 ±$\pm 0.01$ 0.87 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.84 ±$\pm 0.00$ 0.83 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.91 ±$\pm 0.00$
WBC 0.78 ±$\pm 0.01$ 0.73 ±$\pm 0.03$ 0.91 ±$\pm 0.00$ 0.86 ±$\pm 0.00$ 0.73 ±$\pm 0.01$ 0.70 ±$\pm 0.02$ 0.87 ±$\pm 0.00$ 0.80 ±$\pm 0.00$
HB 0.75 ±$\pm 0.01$ 0.74 ±$\pm 0.01$ 0.85 ±$\pm 0.00$ 0.82 ±$\pm 0.00$ 0.74 ±$\pm 0.00$ 0.72 ±$\pm 0.01$ 0.81 ±$\pm 0.00$ 0.78 ±$\pm 0.00$
HCO3 0.79 ±$\pm 0.01$ 0.79 ±$\pm 0.01$ 0.90 ±$\pm 0.00$ 0.90 ±$\pm 0.00$ 0.78 ±$\pm 0.01$ 0.76 ±$\pm 0.00$ 0.88 ±$\pm 0.00$ 0.86 ±$\pm 0.00$
Na 0.80 ±$\pm 0.01$ 0.78 ±$\pm 0.01$ 0.93 ±$\pm 0.00$ 0.92 ±$\pm 0.00$ 0.80 ±$\pm 0.01$ 0.79 ±$\pm 0.01$ 0.91 ±$\pm 0.00$ 0.90 ±$\pm 0.00$
MgSO4 0.81 ±$\pm 0.02$ 0.81 ±$\pm 0.01$ 0.86 ±$\pm 0.00$ 0.83 ±$\pm 0.00$ 0.73 ±$\pm 0.01$ 0.72 ±$\pm 0.01$ 0.81 ±$\pm 0.01$ 0.75 ±$\pm 0.00$
Hep 0.68 ±$\pm 0.01$ 0.65 ±$\pm 0.00$ 0.80 ±$\pm 0.00$ 0.74 ±$\pm 0.00$ 0.65 ±$\pm 0.00$ 0.63 ±$\pm 0.01$ 0.72 ±$\pm 0.00$ 0.66 ±$\pm 0.00$
KCl 0.73 ±$\pm 0.01$ 0.70 ±$\pm 0.00$ 0.80 ±$\pm 0.00$ 0.73 ±$\pm 0.00$ 0.74 ±$\pm 0.01$ 0.66 ±$\pm 0.00$ 0.77 ±$\pm 0.00$ 0.68 ±$\pm 0.00$
NE 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.94 ±$\pm 0.00$
Prop 0.94 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.94 ±$\pm 0.00$
Avg 0.82 ±$\pm 0.09$ 0.80 ±$\pm 0.09$ 0.90 ±$\pm 0.06$ 0.87 ±$\pm 0.08$ 0.80 ±$\pm 0.09$ 0.78 ±$\pm 0.10$ 0.87 ±$\pm 0.08$ 0.83 ±$\pm 0.10$
Table 14: Comparison of VQ-VAE and RQ-VAE on MIMIC-IV and eICU, reporting Column-wise Density Estimation (CDE), Pairwise Column Correlation (PCC), Time Gap, and Event Count metrics.
MIMIC-IV eICU
Metric VQ RQ VQ RQ
CDE $\downarrow$ 0.05 0.04 0.04 0.05
- Categorical 0.05 0.04 0.05 0.06
- Numerical 0.05 0.03 0.03 0.03
- Patientweight 0.28 0.09 0.05 0.03
PCC $\downarrow$ 0.04 0.04 0.07 0.06
Time Gap $\downarrow$ 0.05 0.01 0.03 0.03
# Events $\downarrow$ 0.13 0.02 0.03 0.05
Table 15: Scalability results on MIMIC-IV and eICU for 6-hour and 24-hour observation windows, reporting Column-wise Density Estimation (CDE), Pairwise Column Correlation (PCC), Time Gap, and Event Count metrics.
MIMIC-IV eICU
Obs. size 6h 24h 6h 24h
CDE $\downarrow$ 0.04 0.05 0.06 0.06
PCC $\downarrow$ 0.03 0.02 0.09 0.07
Time Gap $\downarrow$ 0.05 0.07 0.01 0.02
# Events $\downarrow$ 0.07 0.11 0.08 0.01
Table 16: Model Performance Comparison on MIMIC-IV-ED, reporting Column-wise Density Estimation (CDE), Pairwise Column Correlation (PCC), Time Gap, Event Count, and Clinical Utility (AUROC) metrics. The best results are in bold.
Model CDE $\downarrow$ PCC $\downarrow$ Time Gap $\downarrow$ # Events $\downarrow$ Utility $\uparrow$
Real 0.83±$\pm 0.00$
SDV 0.05 0.21 0.46 0.07 0.63±$\pm 0.09$
RC-TGAN 0.55 0.20 0.34 0.06 0.57±$\pm 0.06$
ClavaDDPM 0.16 0.11 0.35 0.06 0.64±$\pm 0.06$
RawMed 0.08 0.02 0.02 0.02 0.79±$\pm 0.05$
F.3 ED dataset generalizability
To assess the generalizability of RawMed across diverse clinical environments, we evaluated its performance on the Emergency Department (ED) dataset from MIMIC-IV-ED, applying the same methodology and evaluation framework as used for the ICU dataset.
Predictive task definition
Due to differences in clinical context, the ICU task definitions were not directly applicable to the ED dataset. Instead, we designed a binary classification task to evaluate the utility of synthetic data, based on four clinically relevant features: heart rate, respiratory rate, morphine administration, and ondansetron administration. The observation window was set to $T=6$ hours, split into two 3-hour segments. The first 3 hours serve as input for the predictive model, with predictions made at the midpoint ($T/2=3$ hours). Binary labels were defined as follows, using the maximum value within the second 3-hour period for heart rate and respiratory rate, and occurrence for medication administration:
  • Heart rate (beats per minute, bpm): 1 if the maximum value is $\geq 120$ bpm (indicating tachycardia, above the normal adult range of 60–100 bpm), 0 otherwise.
  • Respiratory rate (respirations per minute, rpm): 1 if the maximum value is $>24$ rpm (indicating tachypnea, beyond the normal adult range of 12–20 rpm), 0 otherwise.
  • Morphine administration: 1 if administered, 0 if not.
  • Ondansetron administration: 1 if administered, 0 if not.
These thresholds were selected based on established clinical criteria to identify critical conditions in the ED setting.
Results
The evaluation demonstrated that RawMed generalizes effectively to the ED dataset, achieving robust performance across predictive tasks. Table 16 compares RawMed with baseline models (i.e., SDV, RC-TGAN, ClavaDDPM), reporting fidelity metrics (Column-wise Density Estimation, CDE; Pairwise Column Correlation, PCC; Time Gap; Number of Events) and clinical utility. RawMed surpasses baselines in most metrics, with a CDE of 0.08, PCC of 0.02, Time Gap of 0.02, Number of Events error of 0.02, and a Utility score of 0.79, nearing the real data’s Utility of 0.83. Despite differing clinical environments, RawMed’s performance closely matches real data utility and surpasses baselines, indicating potential for generalization to other datasets.
(a) MIMIC-IV / lab
(b) MIMIC-IV / med
(c) MIMIC-IV/ input
(d) eICU / lab
(e) eICU / med
(f) eICU / input
Figure 3: Table-specific visualizations of absolute differences in Pairwise Column Correlation (PCC) matrices between real and synthetic data for MIMIC-IV and eICU tables.
(a) MIMIC-IV / Time Gap (w/ Sim. Events)
(b) eICU / Time Gap (w/ Sim. Events)
(c) MIMIC-IV / Time Gap (w/o Sim. Events)
(d) eICU / Time Gap (w/o Sim. Events)
(e) MIMIC-IV / Absolute Time
(f) eICU / Absolute Time
(g) MIMIC-IV / Event Counts
(h) eICU / Event Counts
Figure 4: Cumulative Distribution Functions (CDFs) of real and synthetic data for MIMIC-IV and eICU datasets, showing Time Gap (with and without simultaneous events), Absolute Time from admission, and Event Counts per patient.

### v2

@@
-Generating Multi-Table Time Series EHR from Latent Space with Minimal Preprocessing
+arXiv:2507.06996v2[cs.LG] 02 Mar 2026
 Abstract
@@
 Unlike previous EHR synthesis methods—which typically generate medical records consisting of expert-chosen features (e.g., a few vital signs, structured codes only)—we introduce RawMed, the first framework to synthesize multi-table, time-series EHR data that closely resembles raw EHRs.
-Using text-based representation and compression techniques, RawMed captures complex structures and temporal dynamics with minimal preprocessing.
+Using text-based representation and compression techniques, RawMed captures complex structures and temporal dynamics with minimal lossy preprocessing.
 We also propose a new evaluation framework for multi-table time-series synthetic EHRs, assessing distributional similarity, inter-table relationships, temporal dynamics, and privacy.
@@
 EHRs systematically store various medical events, such as prescriptions and test results, along with corresponding timestamps in a multi-table relational database, encompassing categorical, numerical, and text data.
-Due to these properties, EHRs serve as a crucial resource in various medical AI studies, including clinical prediction, information retrieval, and question answering Goldstein2017OpportunitiesAC ; Shickel18DeepEHR .
-However, since EHRs contain sensitive personal information, they are subject to privacy regulations that hinder data sharing and research applications Benitez2010EvaluatingRR .
-In response, synthetic data has emerged as a promising solution, offering a privacy-preserving alternative that can still support research and clinical applications yan2022multifaceted ; gonzales2023synthetic .
-Progress in generative models, especially Generative Adversarial Networks (GANs) gan , has opened up new avenues for synthesizing EHR data.
-Early efforts primarily generated discrete features or heterogeneous features, often overlooking the time-series aspect Choi2017GeneratingMD ; torfi2020corgancorrelationcapturingconvolutionalgenerative .
-More recent studies incorporate both time-series and heterogeneous features, gradually approaching a closer reflection of real EHR data emrmgan ; ehrsafe ; flexgenehr ; timediff .
+Due to these properties, EHRs serve as a crucial resource in various medical AI studies, including clinical prediction, information retrieval, and question answering Goldstein et al.
+(2017); Shickel et al.
+(2018).
+However, since EHRs contain sensitive personal information, they are subject to privacy regulations that hinder data sharing and research applications Benitez and Malin (2010).
+In response, synthetic data has emerged as a promising solution, offering a privacy-preserving alternative that can still support research and clinical applications Yan et al.
+(2022); Gonzales et al.
+(2023).
+Progress in generative models, especially Generative Adversarial Networks (GANs) Goodfellow et al.
+(2014), has opened up new avenues for synthesizing EHR data.
+Early efforts primarily generated discrete features or heterogeneous features, often overlooking the time-series aspect Choi et al.
+(2017); Torfi and Fox (2020).
+More recent studies incorporate both time-series and heterogeneous features, gradually approaching a closer reflection of real EHR data Li et al.
+(2023); Yoon et al.
+(2023); He et al.
+(2024); Tian et al.
+(2024).
 Despite this progress, existing EHR synthesis studies still face the following limitations.
 First, there is a high dependence on feature selection.
-Most approaches emrmgan ; ehrsafe ; timediff typically select only a subset of tables and columns of the entire EHR database based on domain knowledge, and generate synthetic data exclusively for the features within those selected columns.
+Most approaches Li et al.
+(2023); Yoon et al.
+(2023); Tian et al.
+(2024) typically select only a subset of tables and columns of the entire EHR database based on domain knowledge, and generate synthetic data exclusively for the features within those selected columns.
 [^1] In this paper, a column denotes a raw data field in an EHR table, while a feature refers to a processed data representation derived from one or more columns.
@@
 For example, if a new research question or analysis requires a feature that was excluded, it is impossible to build predictive models or perform statistical analyses based on that synthetic EHR data.
-Morever, prior research fullfeature1 ; fullfeature2 ; genhpf suggests that models incorporating a broader range of features tend to achieve higher predictive accuracy, implying that richer synthetic data offers wider utility and more substantial value across various downstream tasks.
-Second, many existing approaches rely on complex preprocessing steps.
-Techniques such as numeric binning, one-hot encoding, term normalization, and aggregation are commonly used, but can unintentionally distort the data or obscure meaningful patterns.
+Morever, prior research Rajkomar et al.
+(2018); Nallabasannagari et al.
+(2020); Hur et al.
+(2023) suggests that models incorporating a broader range of features tend to achieve higher predictive accuracy, implying that richer synthetic data offers wider utility and more substantial value across various downstream tasks.
+Second, many existing approaches rely on complex, lossy preprocessing steps.
+Techniques such as numeric binning, term normalization, and aggregation are commonly used, but can unintentionally distort the data or obscure meaningful patterns.
 For instance, aggregating lab values over time may mask sudden anomalies while binning them into discrete ranges may oversimplify subtle trends, reducing the synthetic data’s fidelity for predictive modeling.
@@
 Method [1] [2] [3] [4]
-emrmgan 98 24 $\times$ $\times$
-ehrsafe 90 50 $\times$ $\times$
-flexgenehr 5,373 48 $\times$ $\times$
-timediff 15 276 $\times$ $\times$
+Li et al. (2023) 98 24 $\times$ $\times$
+Yoon et al. (2023) 90 50 $\times$ $\times$
+He et al. (2024) 5,373 48 $\times$ $\times$
+Tian et al. (2024) 15 276 $\times$ $\times$
 RawMed 333,524 243 $\surd$ $\surd$
 To address these challenges, our framework, RawMed, synthesizes raw EHRs, multi-table time-series data that preserve all columns and original values in their database form, as illustrated in Figure 1.
-To implement this approach, RawMed adopts a text-based method, treating EHR data as text to retain original values without additional transformations (e.g., binning, one-hot encoding, normalization).
+To implement this approach, RawMed adopts a text-based method, treating EHR data as text to retain original values without additional transformations (e.g., binning, aggregation, or term normalization).
 This not only reduces the risk of data distortion, but also makes our method generalizable to various EHR data with different database schemas.
 However, as time-series data becomes longer, its textual representation typically grows even longer due to subword tokenization, leading to a significant increase in computational complexity.
-To mitigate this, we employ Residual Quantization rqvae to compress textualized time-series data, enabling autoregressive modeling in a compressed latent space.
+To mitigate this, we employ Residual Quantization Lee et al.
+(2022) to compress textualized time-series data, enabling autoregressive modeling in a compressed latent space.
 This reduces sequence length and computational complexity, allowing RawMed to handle multi-table EHR datasets with numerous columns.
@@
 Early research on EHR synthesis primarily focused on generating single data types (e.g., diagnostic codes) using generative models such as GANs and VAEs.
-Subsequent studies explored mixed data types or incorporated time-series continuity esteban2017real ; biswal2020eva ; zhang2021synteg ; wang2022promptehrconditionalelectronichealthcare ; Theodorou_2023 .
+Subsequent studies explored mixed data types or incorporated time-series continuity Esteban et al.
+(2017); Biswal et al.
+(2020); Zhang et al.
+(2021); Wang and Sun (2022); Theodorou et al.
+(2023b, a).
 However, relatively few studies have addressed both the temporal dynamics and heterogeneous nature of EHR data simultaneously, a gap that has only recently started to gain attention.
-For example, EMR-M-GAN emrmgan pioneered a GAN-based approach for mixed-type time-series EHR data, using a Dual-VAE and Coupled Recurrent Network (CRN) to capture inter-feature correlations and temporal dynamics.
-EHR-Safe ehrsafe also adopted a GAN-based framework designed to handle real-world EHR complexities comprehensively, including missing data, diverse feature types, and both static and time-varying attributes.
-FLEXGEN-EHR flexgenehr integrated static and time-series data, focusing on missing values via optimal transport.
-TIMEDIFF timediff introduced a hybrid diffusion model capable of generating both continuous and discrete time-series data concurrently.
-To the best of our knowledge, no existing studies do not utilize all columns from the original EHR database tables.
-Instead, they rely on a subset of features deemed important by the researchers, and heavily preprocess (e.g., numerical binning, one-hot encoding, normalization, aggregation) the original EHR data to train the generative model.
+For example, EMR-M-GAN Li et al.
+(2023) pioneered a GAN-based approach for mixed-type time-series EHR data, using a Dual-VAE and Coupled Recurrent Network (CRN) to capture inter-feature correlations and temporal dynamics.
+EHR-Safe Yoon et al.
+(2023) also adopted a GAN-based framework designed to handle real-world EHR complexities comprehensively, including missing data, diverse feature types, and both static and time-varying attributes.
+FLEXGEN-EHR He et al.
+(2024) integrated static and time-series data, focusing on missing values via optimal transport.
+TIMEDIFF Tian et al.
+(2024) introduced a hybrid diffusion model capable of generating both continuous and discrete time-series data concurrently.
+To the best of our knowledge, no existing studies utilize all columns from the original EHR database tables.
+Instead, they rely on a subset of features deemed important by the researchers, and heavily preprocess (e.g., numerical binning, aggregation) the original EHR data to train the generative model.
 2.2 Text-based approaches for tabular data generation
-With the rapid progress of NLP, studies have emerged that convert tabular data into textual form for generation purposes great ; tabula .
+With the rapid progress of NLP, studies have emerged that convert tabular data into textual form for generation purposes Borisov et al.
+(2022); Zhao et al.
+(2023).
 This text-based approach offers several advantages: it minimizes complex preprocessing, preserves original values without information loss, and leverages the pre-trained knowledge of language models.
 Such benefits are particularly evident for datasets like EHRs, which often follow varying database schemas and require substantial preprocessing.
-Nevertheless, most existing research on tabular data synthesis focuses on single-table settings great ; tabula , with comparatively few studies addressing relational database realtabformer and no studies synthesizing time-series relational database.
+Nevertheless, most existing research on tabular data synthesis focuses on single-table settings Borisov et al.
+(2022); Zhao et al.
+(2023), with comparatively few studies addressing relational database Solatorio and Dupriez (2023) and no studies synthesizing time-series relational database.
 Moreover, converting tabular data into text increases the length of each row due to subword tokenization, potentially inflating training and inference costs.
@@
 3 Method
-RawMed is a framework for synthesizing multi-table time-series EHRs, capturing complex patient trajectories with minimal preprocessing.
+RawMed is a framework for synthesizing multi-table time-series EHRs, capturing complex patient trajectories with minimal lossy preprocessing.
 Its architecture supports heterogeneous data type and varying event counts using text-based event representation, with two modules: (1) event-level compression for compact event encoding and (2) temporal inter-event modeling for dynamic temporal relationships.
@@
 Although the timestamp is a column in the relational table, we isolate $t_{i}^{p}$ from $a_{i}^{p}$ to highlight its role in temporal ordering.
-[…]
+To create a textual representation, we serialize each event by concatenating the table name $\epsilon_{i}^{p}$, followed by column names and their values, excluding null-valued columns.
+For example, an event from the lab table with $a_{i}^{p}=\{(\text{item},\text{``Glucose''}),\ (\text{value},\ \text{``95''}),\ (\text{uom},\ \text{``mg/dL''})\}$ is serialized as “lab item Glucose value 95 uom mg/dL”.
+Thus, an event is denoted as $e_{i}^{p}=(t_{i}^{p},x_{i}^{p})$, where $x_{i}^{p}$ is the serialized text string.
+Then $x_{i}^{p}$ is tokenized, padded or truncated to a fixed length $L$ (typically 128), and embedded to $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$.
+Given the high dimensionality of these embeddings, compression is essential to model the entire patient sequence $S^{p}$, a challenge also addressed in the context of text-based EHR embeddings Cho et al.
+(2023).
 3.2 Compressing event representations
 To address this, RawMed compress the event text embedding $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, where $L$ is the number of tokens and $F$ is the embedding dimension, into a discrete latent representation $\mathbf{z}_{i}^{p}\in\mathbb{R}^{L_{z}\times F_{z}}$, where $L_{z}$ is the latent sequence length and $F_{z}$ is the latent embedding dimension.
-We adopt neural network-based compression methods, Vector Quantized Variational AutoEncoders (VQ-VAE) vqvae and Residual Quantization (RQ) rqvae , implemented with 1D convolutional neural networks (CNNs).
+We adopt neural network-based compression methods, Vector Quantized Variational AutoEncoders (VQ-VAE) Van Den Oord et al.
+(2017) and Residual Quantization (RQ) Lee et al.
+(2022), implemented with 1D convolutional neural networks (CNNs).
 Encoder and quantization
 The encoder Enc transforms $\mathbf{x}_{i}^{p}$ into $\hat{\mathbf{z}}_{i}^{p}=\texttt{Enc}(\mathbf{x}_{i}^{p})\in\mathbb{R}^{L_{z}\times F_{z}}$, consisting of $L_{z}$ latent vectors.
-Each latent vector, denoted $\hat{\mathbf{z}}\in\mathbb{R}^{F_{z}}$ for simplicity, is quantized to $\mathbf{z}\in\mathbb{R}^{F_{z}}$ by assigning it to the nearest entry in the codebook $C=\{(k,\texttt{lut}(k))\mid k=1,\dots,K\}$, where $\texttt{lut}(k)\in\mathbb{R}^{F_{z}}$ is the embedding for index $k$ [^2] The function lut refers to a look-up table that maps indices to their corresponding embeddings.
+Each latent vector, denoted $\hat{\mathbf{z}}\in\mathbb{R}^{F_{z}}$ for simplicity, is quantized to $\mathbf{z}\in\mathbb{R}^{F_{z}}$ by assigning it to the nearest entry in the codebook $C=\{(k,\texttt{lut}(k))\mid k=1,\dots,K\}$, where $\texttt{lut}(k)\in\mathbb{R}^{F_{z}}$ is the embedding for index kk.
+[^2] The function lut refers to a look-up table that maps indices to their corresponding embeddings.
 The quantization is defined as:
@@
 $\texttt{RQ}(\hat{\mathbf{z}};C,D)=(k_{1},\dots,k_{D})\in[K]^{D},\quad\mathbf{z}=\sum_{d=1}^{D}\texttt{lut}(k_{d}),$
-where $D$ is the quantization depth (i.e., $D$=1 for VQ).
+where $D$ is the quantization depth (i.e., DD=1 for VQ).
 The RQ process initializes the residual vector as $\mathbf{r}_{0}=\hat{\mathbf{z}}$.
@@
 Next, the event tokens are mapped to latent representations $\tilde{\mathbf{z}}_{i}^{p}$ via codebook $C$, decoded by Dec into text embeddings $\tilde{\mathbf{x}}_{i}^{p}$, and converted to text (e.g., lab item Glucose value 95 uom mg/dL).
-Time tokens are detokenized into timestamps (e.g., $[7,2]\to 720~{}\text{minutes}$).
+Time tokens are detokenized into timestamps (e.g., $7,2]\to 720~\text{minutes}$).
 These text-based events and timestamps are then converted to relational tables.
@@
 Dataset
-In this study, we used two publicly available EHR datasets, MIMIC-IV mimiciv and eICU eicu .
-Adhering to the GenHPF genhpf framework, which utilizes all EHR features for diverse prediction tasks, we focused on patients aged 18 years or older who had been admitted to the ICU for more than 12 hours.
+In this study, we used two publicly available EHR datasets, MIMIC-IV Johnson et al.
+(2020) and eICU Pollard et al.
+(2018).
+Adhering to the GenHPF Hur et al.
+(2023) framework, which utilizes all EHR features for diverse prediction tasks, we focused on patients aged 18 years or older who had been admitted to the ICU for more than 12 hours.
 Based on this framework, we integrated three time-series tables (laboratory tests, prescriptions, and input events) from both datasets.
-To assess the generalizability of our approach across diverse clinical settings, we conducted additional experiments on the MIMIC-IV-ED mimicived emergency department dataset, with results reported in Appendix F.3.
+To assess the generalizability of our approach across diverse clinical settings, we conducted additional experiments on the MIMIC-IV-ED Johnson et al.
+(2023) emergency department dataset, with results reported in Appendix F.3.
 After preprocessing, the datasets comprised 62,704 patients (about 7 million rows) from MIMIC-IV and 143,812 (about 7 million rows) from eICU, with summary statistics detailed in Table 7.
@@
 Because existing methods rely on predetermined feature selection and complex preprocessing steps, it is infeasible to extend them to the broader generative scope pursued in this work (i.e., all elements in the original database structure).
-To provide a baseline, we adapt RealTabFormer realtabformer , designed for text-based relational databases rather than time-series relational databases.
+To provide a baseline, we adapt RealTabFormer Solatorio and Dupriez (2023), designed for text-based relational databases rather than time-series relational databases.
 Built on GPT-2, RealTabFormer is constrained by limited context length, restricting the amount of data it can process.
-To address this, we concatenate time-series data in chronological order and apply QLoRA dettmers2024qlora to Llama 3.1dubey2024llama , rather than GPT-2 gpt2 , allowing for longer sequences.
+To address this, we concatenate time-series data in chronological order and apply QLoRA Dettmers et al.
+(2024) to Llama 3.1Dubey et al.
+(2024), rather than GPT-2 Radford et al.
+(2019), allowing for longer sequences.
 Moreover, RealTabFormer does not compress data but instead models it directly, effectively serving as a no-compression baseline.
-Additionally, we include baselines using SDV sdv (Gaussian Copula-based), RC-TGAN rctgan (GAN-based), and ClavaDDPM clavaddpm (diffusion-based) for multi-table generation.
+Additionally, we include baselines using SDV Patki et al.
+(2016) (Gaussian Copula-based), RC-TGAN Gueye et al.
+(2023) (GAN-based), and ClavaDDPM Pang et al.
+(2024) (diffusion-based) for multi-table generation.
 These methods, while not designed for time-series multi-table data, can generate synthetic data by modeling timestamps as standard columns within each table, as is typical in EHR database structures.
@@
 Evaluation of synthetic multi-table time-series data, particularly for raw EHRs, remains underexplored.
-Existing frameworks, such as SDMetrics [^3] https://docs.sdv.dev/sdmetrics and Synthcity synthcity , primarily focus on single-table evaluation.
+Existing frameworks, such as SDMetrics [^3] https://docs.sdv.dev/sdmetrics and Synthcity Qian et al.
+(2023), primarily focus on single-table evaluation.
 In multi-table scenarios, these frameworks typically average metrics across individual tables or merge tables into a single one, obscuring inter-table relationships and temporal dynamics critical to EHR data.
@@
 Pairwise Column Correlation (PCC) and Item-specific Pairwise Column Correlation (I-PCC) assess inter-column dependencies in synthetic data by comparing correlation matrices of real and synthetic tables.
-For PCC, matrices are computed using Pearson’s correlation coefficient (range: $[-1,1]$) for numeric-numeric pairs, Theil’s U statistic (range: $[0,1]$) for categorical-categorical pairs, and correlation ratio (range: $[0,1]$) for categorical-numeric pairs.
+For PCC, matrices are computed using Pearson’s correlation coefficient (range: [1,1][-1,1]) for numeric-numeric pairs, Theil’s U statistic (range: $[0,1]$) for categorical-categorical pairs, and correlation ratio (range: $[0,1]$) for categorical-numeric pairs.
 The mean absolute difference ($\mu_{\mathrm{abs}}$) quantifies dependency fidelity by averaging element-wise matrix differences.
@@
 Unlike single-table data, multi-table EHR data exhibit significant variability in event types and lengths per patient, necessitating comprehensive representations.
-To address this variability, we employ two methods: GenHPF genhpf , which lists all events sequentially as a single textual sequence, and MEDS-TAB medstab , which aggregates events into fixed time intervals (see Appendix E.2.1 for details).
+To address this variability, we employ two methods: GenHPF Hur et al.
+(2023), which lists all events sequentially as a single textual sequence, and MEDS-TAB Oufattole et al.
+(2024), which aggregates events into fixed time intervals (see Appendix E.2.1 for details).
 For evaluation, we define 11 clinical prediction tasks (see Table 10), and report the Area Under the Receiver Operating Characteristic Curve (AUROC) for models trained on synthetic data compared to those trained on real data, using real test data.
-We also perform Membership Inference Attacks (MIA) mia to evaluate privacy leakage in synthetic data by measuring distances between synthetic samples and the training/test dataset to infer membership, where successful identification indicates inadequate privacy protection.
+We also perform Membership Inference Attacks (MIA) Shokri et al.
+(2017) to evaluate privacy leakage in synthetic data by measuring distances between synthetic samples and the training/test dataset to infer membership, where successful identification indicates inadequate privacy protection.
 To evaluate the temporal fidelity, we introduce two metrics: Time Gap and Event Count.
@@
 Event Count assesses the distribution of event counts per patient, also using the KS statistic, to evaluate whether synthetic data accurately reflects the frequency of clinical events.
-Finally, Next Event Prediction evaluates temporal sequence dynamics using an LSTM lstm -based model to predict the next event’s item or drug name, formulated as a multi-label classification task for concurrent events.
+Finally, Next Event Prediction evaluates temporal sequence dynamics using an LSTM Hochreiter and Schmidhuber (1997)-based model to predict the next event’s item or drug name, formulated as a multi-label classification task for concurrent events.
 The model is trained separately on synthetic and real data, with performance evaluated on real test data using the F1 score, quantifying the synthetic data’s ability to capture event sequence patterns.
@@
 RC-TGAN $0.51\text{\scriptsize$\pm 0.14$}$ $0.51\text{\scriptsize$\pm 0.13$}$ $0.47\text{\scriptsize$\pm 0.13$}$ $0.48\text{\scriptsize$\pm 0.11$}$
-ClavaDDPM $0.68\text{\scriptsize$\pm 0.19$}$ $0.64\text{\scriptsize$\pm 0.17$}$ $0.66\text{\scriptsize$\pm 0.14$}$ $0.63\text{\scriptsize$\pm 0.12$}$
+ClavaDDPM $0.68\text{\scriptsize$\pm 0.19$}$ $4\text{\scriptsize$\pm 0.17$} 0.66±0$$6\text{\scriptsize$\pm 0.14$}$ $0.63\text{\scriptsize$\pm 0.12$}$
 RawMed $\mathbf{0.87}\text{\scriptsize$\pm 0.08$}$ $\mathbf{0.80}\text{\scriptsize$\pm 0.09$}$ $\mathbf{0.83}\text{\scriptsize$\pm 0.10$}$ $\mathbf{0.78}\text{\scriptsize$\pm 0.10$}$
@@
 Model MIMIC-IV eICU MIMIC-IV eICU MIMIC-IV eICU MIMIC-IV eICU
-Real $0.18\text{\scriptsize$\pm 0.06$}$ $0.30\text{\scriptsize$\pm 0.00$}$
+Real $0.18\text{\scriptsize$\pm 0.06$}$ $0.30\text{\scriptsize$\pm 0.00$}$ - - - - - -
 SDV $0.05\text{\scriptsize$\pm 0.01$}$ $0.13\text{\scriptsize$\pm 0.00$}$ 0.76 0.50 0.46 0.36 0.499 0.502
-RC-TGAN $0.02\text{\scriptsize$\pm 0.00$}$ $0.07\text{\scriptsize$\pm 0.01$}$ 0.43 0.39 0.04 0.03 0.500 0.500
+RC-TGAN $02\text{\scriptsize$\pm 0.00$} 0.0$$07\text{\scriptsize$\pm 0.01$}$ 0.43 0.39 0.04 0.03 0.500 0.500
 ClavaDDPM $0.06\text{\scriptsize$\pm 0.00$}$ $0.12\text{\scriptsize$\pm 0.00$}$ 0.48 0.41 0.11 0.05 0.500 0.499
@@
 Compressed vs. Non-compressed approaches
-We compared RealTabFormer (RTF, non-compressed) and RawMed (compressed) to evaluate the effect of sequence compression on synthetic data quality.444 Experiments typically used a 12-hour observation window for training and synthesis, but both RTF and RawMed used a 6-hour window in this experiment due to RTF’s context-size limits.
+We compared RealTabFormer (RTF, non-compressed) and RawMed (compressed) to evaluate the effect of sequence compression on synthetic data quality.
+[^4] Experiments typically used a 12-hour observation window for training and synthesis, but both RTF and RawMed used a 6-hour window in this experiment due to RTF’s context-size limits.
 We synthesized about 20k samples to match RTF-generated sample counts, as generating the original dataset size was infeasible for RTF.
@@
 In subsequent work, we plan to expand the range of tables, implement conditional generation, and more fully integrate static features into the model design, thereby providing more realistic synthetic EHR data for a broader array of medical AI tasks.
+Acknowledgments and Disclosure of Funding
+This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants (No.RS-2019-II190075), the Korea Health Industry Development Institute (KHIDI) grant (No.HR21C0198, No.RS-2025-02213750) and the National Research Foundation of Korea (NRF) grants (NRF-2020H1D3A2A03100945), funded by the Korea government (MSIT, MOHW).
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and introduction accurately state RawMed’s contributions, including multi-table EHR generation and a novel evaluation framework.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Section 7 discusses limitations and future work to address these constraints.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper focuses on empirical results without theoretical theorems or proofs. Methodological details are provided in Section 3.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Section 4 provides details on datasets, baselines, and evaluation metrics, with further implementation details in Appendix E.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Supplementary material includes source code, utilizing the open-source MIMIC-IV and eICU datasets.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Appendix E and Section 4 provide comprehensive details on data splits, hyperparameters, and training procedures.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Tables 4, 11, 12 and 13 provide mean and standard deviation across multiple runs.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Appendix E specifies GPU types and training times for experiments.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We thoroughly reviewed the NeurIPS Code of Ethics and ensured full compliance with its guidelines.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Section 7 discusses the broader impacts of our research.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper uses public datasets (MIMIC-IV, eICU) and generates synthetic EHR data with privacy protections evaluated via Membership Inference Attacks (MIA, Table 4).
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The paper cites MIMIC-IV and eICU datasets with their respective licenses.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The evaluation framework and synthetic data generation code are documented with public access details.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing or human subject research.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper uses de-identified publicly available datasets, requiring no IRB approval.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: The paper uses Meta-LLaMA-3.1-8B for RealTabFormer, described in Appendix E as a baseline method.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Feature definitions in EHR data
@@
 The Possible Feature Count assumes all possible combinations of categorical values exist for each item, providing a theoretical upper bound. It is calculated as:
-Possible Feature $\text{Possible Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg{(}\sum_{i\in I_{\epsilon}}\Big{(}\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|\Big{)}\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg{)},$
+$\text{Possible Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg(\sum_{i\in I_{\epsilon}}\Big(\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|\Big)\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg),$
 where $\prod_{c\in C_{\epsilon,\text{cat}}}|V_{i,c}|$ represents the number of possible categorical combinations, and $|C_{\epsilon,\text{num}}|+1$ accounts for the numerical columns plus a indicator feature representing the presence or absence of the categorical combination.
@@
 The Actual Feature Count considers only the categorical combinations observed in the data, reflecting dependencies or sparsity that reduce the number of combinations. It is calculated as:
-Actual Feature $\text{Actual Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg{(}\sum_{i\in I_{\epsilon}}\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg{)},$
+$\text{Actual Feature Count}=\sum_{\epsilon\in\mathcal{E}}\Bigg(\sum_{i\in I_{\epsilon}}\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|\times\left(|C_{\epsilon,\text{num}}|+1\right)\Bigg),$
 where $\left|C_{\epsilon,\text{cat}}^{\text{unique}}(i)\right|$ is the number of actual combinations, and $|C_{\epsilon,\text{num}}|+1$ accounts for numerical columns and the presence/absence feature.
@@
 The $\mu_{\mathrm{abs}}$ is calculated for each item’s matrix and averaged across items.
-As with I-CDE, items with fewer than 0.1% of records are excluded.
+Items with fewer than 100 records are excluded.
     • Justification: I-PCC evaluates item-specific inter-column dependencies, ensuring that synthetic data captures relationships within subsets of EHR data (e.g., correlations between dosage and frequency for a specific drug).
@@
   • Clinical Utility: To assess clinical utility, we evaluate the performance of synthetic data on downstream predictive tasks.
-We defined 11 clinical prediction tasks (e.g., predicting creatinine and hemoglobin levels), following the task definitions from genhpf .
+We defined 11 clinical prediction tasks (e.g., predicting creatinine and hemoglobin levels), following the task definitions from Hur et al.
+(2023).
 Details are provided in Table 10.
@@
   • Next Event Prediction: Next Event Prediction evaluates the synthetic data’s ability to capture temporal sequence dynamics.
-An LSTM-based model (lstm, ) predicts the next event’s item or drug name, formulated as a multi-label classification task to handle concurrent events.
+An LSTM-based model (Hochreiter and Schmidhuber, 1997) predicts the next event’s item or drug name, formulated as a multi-label classification task to handle concurrent events.
 The model is trained on synthetic data and evaluated on real test data using the F1 score, which balances precision and recall for multi-label predictions.
@@
 This section details the architectural components of the RawMed framework, consisting of two distinct modules: event compression and inter-event temporal modeling, providing specifics beyond the main text.
-The event compression module transforms serialized clinical event text into compact latent representations, leveraging a Vector Quantized Variational AutoEncoder (VQ-VAE) vqvae or Residual Quantization (RQ) framework rqvae , both implemented using 1D convolutional neural networks (CNNs).
+The event compression module transforms serialized clinical event text into compact latent representations, leveraging a Vector Quantized Variational AutoEncoder (VQ-VAE) Van Den Oord et al.
+(2017) or Residual Quantization (RQ) framework Lee et al.
+(2022), both implemented using 1D convolutional neural networks (CNNs).
 The processing pipeline begins with the construction of input embeddings, followed by encoding, quantization, and decoding stages.
 The input embedding, denoted $\mathbf{x}_{i}^{p}\in\mathbb{R}^{L\times F}$, integrates three components adapted from an established text-based EHR prediction framework, GenHPF, to ensure robust representation of clinical data.
-The textual embedding ($\mathbf{x}_{\text{text},i}^{p}$) employs framework-derived embeddings for serialized event text, such as “lab item Glucose value 95 uom mg/dL”, tokenized using a Bio+Clinical BERT tokenizer bioclinicalbert .
+The textual embedding ($\mathbf{x}_{\text{text},i}^{p}$) employs framework-derived embeddings for serialized event text, such as “lab item Glucose value 95 uom mg/dL”, tokenized using a Bio+Clinical BERT tokenizer Alsentzer et al.
+(2019).
 The type embedding ($\mathbf{x}_{\text{type},i}^{p}$) assigns categorical labels to tokens, distinguishing table names, column names, and values to capture the relational structure of the data.
@@
 D.2 Postprocessing
-This appendix elaborates on the postprocessing steps applied to convert synthetic patient trajectories into relational tables, aligning with the structural and semantic consistency of real data.
-The core steps of event-level verification and patient-level validation, which ensure valid event formats and temporal consistency, are implemented in Algorithm 1.
-Here, we detail the additional postprocessing steps that enforce column-specific constraints, as referenced in Section 3.4, to refine the synthetic tables’ statistical fidelity.
+This appendix elaborates on the two-stage postprocessing pipeline applied to convert synthetic patient trajectories into relational tables, aligning with the structural and semantic consistency of real data.
+The first stage, converting text to tabular format, involves the core steps of event-level verification and patient-level validation (Algorithm 1).
+The second stage, enhancing data quality, applies additional column-specific constraints to refine the synthetic tables’ statistical fidelity, as referenced in Section 3.4.
+Column-Specific Constraint Enforcement
 After validated sequences are converted into relational tables, further postprocessing ensures compliance with column-specific constraints, addressing potential errors from the generative process (e.g., quantization or autoregressive sampling).
@@
 These steps ensure that the final relational tables adhere to the structural and statistical characteristics of the real data, removing outliers and ensuring both numerical and categorical validity.
-Input: Text-based events $E$, timestamps $\mathcal{T}$, valid column names $\mathcal{C}$, valid table names $\mathcal{E}$, observation window $\mathcal{T}_{\text{max}}$
-Output: Relational tables $\mathcal{R}$
-Initialize $\mathcal{R}\leftarrow\emptyset$ (dictionary of tables)
-for each event $e_{i}\in E$ with timestamp $t_{i}\in\mathcal{T}$ do
-       Parse $e_{i}$ into table name $\tau$ and column-value pairs $\{(c_{i,j},v_{i,j})\}$
-      if $\tau\notin\mathcal{E}$ then
-             Discard $e_{i}$
-            Continue
-       end if
-      if $e_{i}$ does not conform to column-value pair format then
-             Discard $e_{i}$
-            Continue
-       end if
-      for each $(c_{i,j},v_{i,j})$ in $\{(c_{i,j},v_{i,j})\}$ do
-             if $c_{i,j}\notin\mathcal{C}$ then
-                   $c_{i,j}\leftarrow\arg\min_{c\in\mathcal{C}}\text{Levenshtein}(c_{i,j},c)$
-             end if
-            if $c_{i,j}$ is numeric and $v_{i,j}$ contains non-numeric characters then
-                   $v_{i,j}\leftarrow\text{RemoveNonNumeric}(v_{i,j})$
-             end if
-       end for
-      Add $(t_{i},\{(c_{i,j},v_{i,j})\})$ to $\mathcal{R}[\tau]$
-end for
-for each table $\mathcal{R}[\tau]$ in $\mathcal{R}$ do
-       Sort events by timestamp $t_{i}$ in ascending order
-      for each event $(t_{i},\{(c_{i,j},v_{i,j})\})$ in $\mathcal{R}[\tau]$ do
-             if $i>1$ and $t_{i}<t_{i-1}$ or $t_{i}>\mathcal{T}_{\text{max}}$ then
-                   Discard $(t_{i},\{(c_{i,j},v_{i,j})\})$ and all events $(t_{k},\{(c_{k,j},v_{k,j})\})$ for $k\geq i$
-                  Break
-             end if
-       end for
-end for
-return $\mathcal{R}$
-Algorithm 1 Postprocessing for Relational Table Construction
+Modification and Rejection Rate Analysis
+To quantify the impact of our validation protocols, we analyzed the modification and rejection rates across both postprocessing stages for the eICU and MIMIC-IV datasets.
+During the initial stage, converting text to tabular format, $0.42\%$ of events were modified and $0.07\%$ were rejected in eICU.
+For MIMIC-IV, these rates were $0.73\%$ and $0.17\%$, respectively.
+At the patient level, however, rejection rates were notably higher, reaching $2.71\%$ for eICU and $18.23\%$ for MIMIC-IV.
+In the subsequent stage, enhancing data quality, stricter column-specific constraints were enforced.
+This led to event-level rates of $1.39\%$ modified and $0.49\%$ rejected for eICU, and $1.07\%$ modified and $0.79\%$ rejected for MIMIC-IV.
+These stringent checks led to a substantial increase in patient-level rejection rates, which rose to $16.90\%$ in eICU and $54.53\%$ in MIMIC-IV.
+The pronounced disparity between patient-level and event-level rejection rates is a direct consequence of our stringent data integrity policy.
+A single erroneous event necessitates the rejection of the entire patient trajectory.
+This effect is particularly notable for datasets with long sequences, (e.g., with hundreds of events), all of which must be accurate for the trajectory to be retained.
+We emphasize that this full post-processing is crucial for RawMed’s data integrity.
+We do not recommend partial post-processing, as it fundamentally compromises the quality of the synthetic data.
 Appendix E Training details and hyperparameters
@@
     • Generation: Data generation uses top-$k$ sampling with $k=250$ for MIMIC-IV and $k=150$ for eICU.
-  • RealTabFormer realtabformer : Text-based autoregressive model for generating relational database.
+  • RealTabFormer Solatorio and Dupriez (2023): Text-based autoregressive model for generating relational database.
 This model involves fine-tuning the Meta-LLaMA-3.1-8B model and a generation step using top-p sampling.
-    • Fine-tuning: The Meta-LLaMA-3.1-8B model dubey2024llama was fine-tuned using QLoRA dettmers2024qlora (rank $r$=128, LoRA scaling factor $\alpha$=128, LoRA dropout=0.05) and Flash Attention-2 for memory optimization.
+    • Fine-tuning: The Meta-LLaMA-3.1-8B model Dubey et al.
+(2024) was fine-tuned using QLoRA Dettmers et al.
+(2024) (rank rr=128, LoRA scaling factor $\alpha$=128, LoRA dropout=0.05) and Flash Attention-2 for memory optimization.
 Training used a maximum sequence length of 11,240 tokens, matching the input data.
@@
     • Generation: Used the fine-tuned Meta-LLaMA-3.1-8B model with top-p sampling, setting the temperature to 1 and a threshold of 0.7.
-  • SDV sdv : The Synthetic Data Vault (SDV) uses the Hierarchical ML Algorithm (HMA) Synthesizer to generate synthetic relational data, modeling individual tables with Gaussian Copulas while preserving parent-child relationships.The SDV framework’s default hyperparameters are applied for training.
-  • RC-TGANrctgan A GAN-based model designed to generate synthetic relational databases by modeling parent-child relationships.
+  • SDV Patki et al.
+(2016): The Synthetic Data Vault (SDV) uses the Hierarchical ML Algorithm (HMA) Synthesizer to generate synthetic relational data, modeling individual tables with Gaussian Copulas while preserving parent-child relationships.The SDV framework’s default hyperparameters are applied for training.
+  • RC-TGANGueye et al. (2023) A GAN-based model designed to generate synthetic relational databases by modeling parent-child relationships.
     • Training and generation: Uses default hyperparameters from the official repository, with batch size set to 500,000, the maximum feasible value for the NVIDIA RTX 3090 GPU (default: 500).
@@
 Training took less than 10 hours on a single NVIDIA RTX 3090 GPU without early stopping.
-  • ClavaDDPMclavaddpm A diffusion-based model for multi-table data generation, Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM) leverage Denoising Diffusion Probabilistic Models (DDPMs) to model complex tabular data distributions.
+  • ClavaDDPMPang et al.
+(2024) A diffusion-based model for multi-table data generation, Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM) leverage Denoising Diffusion Probabilistic Models (DDPMs) to model complex tabular data distributions.
 By using clustering labels as intermediaries, ClavaDDPM captures long-range dependencies across interconnected tables, particularly through foreign key constraints.
@@
 Training was completed in under 3 hours on a single NVIDIA A6000 GPU.
+Input: Set of patients $\mathcal{P}$, Text-based events $E$, timestamps $\mathcal{T}$, valid column names $\mathcal{C}$, valid table names $\mathcal{E}$, observation window $\mathcal{T}_{\text{max}}$
+Output: Relational tables $\mathcal{R}$
+Initialize $\mathcal{R}\leftarrow\text{dict of }\tau\to\text{list}$ (a dictionary mapping table names to lists of event records)
+Initialize $\mathcal{P}_{\text{valid}}\leftarrow\emptyset$
+for each patient $p\in\mathcal{P}$ do
+   Initialize $\mathcal{R}^{\prime}\leftarrow\text{dict of }\tau\to\text{list}$ (temporary tables for patient $p$)
+   Initialize $\text{valid\_event\_count}\leftarrow 0$
+ for each event $e_{i}^{p}\in E^{p}$ with timestamp $t_{i}^{p}\in\mathcal{T}^{p}$ do
+      Parse $e_{i}^{p}$ into table name $\tau$ and column-value pairs $\{(c_{i,j}^{p},v_{i,j}^{p})\}$
+    if $\tau\notin\mathcal{E}$ then
+         Continue
+      end if
+    if $e_{i}^{p}$ does not conform to column-value pair format then
+         Continue
+      end if
+    for each $(c_{i,j}^{p},v_{i,j^{p}})$ in $\{(c_{i,j}^{p},v_{i,j}^{p})\}$ do
+       if $c_{i,j}^{p}\notin\mathcal{C}$ then
+          $c_{i,j}^{p}\leftarrow\arg\min_{c\in\mathcal{C}}\text{Levenshtein}(c_{i,j}^{p},c)$
+         end if
+       if $c_{i,j}^{p}$ is numeric column and $v_{i,j}^{p}$ contains non-numeric characters then
+          $v_{i,j}^{p}\leftarrow\text{RemoveNonNumeric}(v_{i,j}^{p})$
+          if $v_{i,j}^{p}$ is not numeric then
+               Continue
+            end if
+         end if
+      end for
+     Add $(t_{i}^{p},\{(c_{i,j}^{p},v_{i,j}^{p})\})$ to $\mathcal{R}^{\prime}[\tau]$
+    $\text{valid\_event\_count}\leftarrow\text{valid\_event\_count}+1$
+   end for
+ if valid_event_count == len($E^{p}$) then
+      Add $p$ to $\mathcal{P}_{\text{valid}}$
+     Add $\mathcal{R}^{\prime}[\tau]$ to $\mathcal{R}[\tau]$
+   end if
+end for
+Sort $\mathcal{R}$ by timestamp in ascending order
+for each patient $p$ in $\mathcal{P}_{\text{valid}}$ do
+ for each event $(t_{i}^{p},\{(c_{i,j}^{p},v_{i,j}^{p})\})$ in $\mathcal{R}$ do
+    if $i>1$ and $t_{i}^{p}<t_{i-1}^{p}$ or $t_{i}^{p}>\mathcal{T}_{\text{max}}$ then
+         Discard $(t_{i}^{p},\{(c_{i,j}^{p},v_{i,j}^{p})\})$ and all events $(t_{k}^{p},\{(c_{k,j}^{p},v_{k,j}^{p})\})$ for $k\geq i$ from $\mathcal{R}$
+        Break
+      end if
+   end for
+end for
+return $\mathcal{R}$
+Algorithm 1 Postprocessing for Relational Table Construction
 E.2 Predictive models for evaluation
@@
 E.2.1 Clinical utility
-  • GenHPF genhpf : GenHPF addresses the heterogeneity of multi-table EHR data by transforming all patient events into a single hierarchical textual sequence.
+  • GenHPF Hur et al.
+(2023): GenHPF addresses the heterogeneity of multi-table EHR data by transforming all patient events into a single hierarchical textual sequence.
 This method sequentially lists events such as medications, infusions, and lab results in chronological order, requiring minimal preprocessing.
@@
 Training occurred on a single NVIDIA A6000 GPU, targeting multi-task prediction for EHR lab tests and medications, with early stopping after 10 epochs of no improvement.
-  • MEDS-TAB medstab : MEDS-TAB is an automated tabularization tool that converts Medical Event Data Standard (MEDS)-formatted EHR data into standardized tabular representations by aggregating events into fixed time intervals.
+  • MEDS-TAB Oufattole et al.
+(2024): MEDS-TAB is an automated tabularization tool that converts Medical Event Data Standard (MEDS)-formatted EHR data into standardized tabular representations by aggregating events into fixed time intervals.
 It processes irregularly sampled time-series data, supporting various aggregation functions (e.g., sum, count, mean) and window sizes to handle diverse data types, including static codes, numerical values, and time-series events.
@@
 Col. Name Type CDE $\downarrow$ I-CDE $\downarrow$ Real Null (%) Syn Null (%) TRTR TSTR
-L/itemid C 0.06 - 0.00 0.00 66.08 ±$\pm 24.72$ 78.87 ±$\pm 23.05$
-L/value C 0.07 0.14 5.78 4.51 0.00 ±$\pm 0.00$ 0.77 ±$\pm 0.02$
-L/valueuom C 0.04 0.03 13.78 12.49 0.07 ±$\pm 0.00$ 2.50 ±$\pm 0.02$
+L/itemid C 0.06 - 0.00 0.00 66.08 $\pm 24.72$ 78.87 $\pm 23.05$
+L/value C 0.07 0.14 5.78 4.51 0.00 $\pm 0.00$ 0.77 $\pm 0.02$
+L/valueuom C 0.04 0.03 13.78 12.49 0.07 $\pm 0.00$ 2.50 $\pm 0.02$
 L/flag C 0.00 0.00 62.27 64.18 - -
-L/priority C 0.03 0.04 27.34 29.48 44.16 ±$\pm 0.08$ 44.32 ±$\pm 0.19$
-L/comments C 0.11 0.22 83.41 86.15 57.34 ±$\pm 0.61$ 63.52 ±$\pm 2.04$
-L/valuenum N 0.01 0.06 11.13 9.46 116.64 ±$\pm 0.76$ 86.37 ±$\pm 1.99$
-L/ref_range_lower N 0.02 0.01 19.04 17.15 60.86 ±$\pm 0.21$ 63.10 ±$\pm 0.20$
-L/ref_range_upper N 0.01 0.01 19.04 17.15 30.45 ±$\pm 0.66$ 30.01 ±$\pm 0.36$
-I/itemid C 0.09 - 0.00 0.00 28.17 ±$\pm 3.03$ 29.19 ±$\pm 0.13$
-I/amountuom C 0.02 0.02 0.00 0.00 0.07 ±$\pm 0.00$ 0.13 ±$\pm 0.00$
-I/rateuom C 0.05 0.05 37.49 39.54 0.02 ±$\pm 0.00$ 0.07 ±$\pm 0.00$
-I/ordercategoryname C 0.03 0.03 0.00 0.00 0.16 ±$\pm 0.00$ 0.19 ±$\pm 0.00$
-I/secondaryordercategoryname C 0.02 0.01 25.82 26.57 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
-I/ordercomponenttypedescription C 0.01 0.01 0.00 0.00 0.00 ±$\pm 0.00$ 0.01 ±$\pm 0.00$
-I/ordercategorydescription C 0.02 0.03 0.00 0.00 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
+L/priority C 0.03 0.04 27.34 29.48 44.16 $\pm 0.08$ 44.32 $\pm 0.19$
+L/comments C 0.11 0.22 83.41 86.15 57.34 $\pm 0.61$ 63.52 $\pm 2.04$
+L/valuenum N 0.01 0.06 11.13 9.46 116.64 $\pm 0.76$ 86.37 $\pm 1.99$
+L/ref_range_lower N 0.02 0.01 19.04 17.15 60.86 $\pm 0.21$ 63.10 $\pm 0.20$
+L/ref_range_upper N 0.01 0.01 19.04 17.15 30.45 $\pm 0.66$ 30.01 $\pm 0.36$
+I/itemid C 0.09 - 0.00 0.00 28.17 $\pm 3.03$ 29.19 $\pm 0.13$
+I/amountuom C 0.02 0.02 0.00 0.00 0.07 $\pm 0.00$ 0.13 $\pm 0.00$
+I/rateuom C 0.05 0.05 37.49 39.54 0.02 $\pm 0.00$ 0.07 $\pm 0.00$
+I/ordercategoryname C 0.03 0.03 0.00 0.00 0.16 $\pm 0.00$ 0.19 $\pm 0.00$
+I/secondaryordercategoryname C 0.02 0.01 25.82 26.57 0.00 $\pm 0.00$ 0.00 $\pm 0.00$
+I/ordercomponenttypedescription C 0.01 0.01 0.00 0.00 0.00 $\pm 0.00$ 0.01 $\pm 0.00$
+I/ordercategorydescription C 0.02 0.03 0.00 0.00 0.00 $\pm 0.00$ 0.00 $\pm 0.00$
 I/totalamountuom C 0.00 0.00 12.35 13.59 - -
-I/isopenbag C 0.01 0.01 0.00 0.00 0.00 ±$\pm 0.00$ 0.00 ±$\pm 0.00$
-I/amount N 0.03 0.08 0.00 0.05 88.26 ±$\pm 0.87$ 87.75 ±$\pm 2.87$
-I/rate N 0.04 0.13 37.49 39.54 36.47 ±$\pm 1.69$ 62.34 ±$\pm 1.11$
-I/patientweight N 0.09 - 0.00 0.00 21.23 ±$\pm 0.01$ 21.78 ±$\pm 0.02$
-I/totalamount N 0.02 0.07 12.38 13.63 6.77 ±$\pm 0.08$ 8.81 ±$\pm 0.02$
-I/originalamount N 0.03 0.12 0.17 0.10 28.00 ±$\pm 1.23$ 30.78 ±$\pm 0.34$
-I/originalrate N 0.03 0.09 5.83 4.46 41.72 ±$\pm 3.88$ 48.13 ±$\pm 2.02$
-M/drug C 0.08 - 0.00 0.00 31.00 ±$\pm 0.38$ 40.53 ±$\pm 1.01$
-M/drug_type C 0.00 0.00 0.00 0.00 0.82 ±$\pm 0.07$ 1.41 ±$\pm 0.04$
-M/prod_strength C 0.08 0.05 0.09 0.04 17.69 ±$\pm 0.63$ 37.79 ±$\pm 0.43$
-M/dose_unit_rx C 0.03 0.02 0.09 0.04 5.93 ±$\pm 0.11$ 13.06 ±$\pm 0.28$
-M/form_unit_disp C 0.03 0.04 0.09 0.05 8.16 ±$\pm 0.21$ 21.07 ±$\pm 0.45$
-M/route C 0.04 0.04 0.09 0.04 31.99 ±$\pm 0.49$ 40.59 ±$\pm 0.30$
-M/doses_per_24_hrs N 0.02 0.06 59.97 61.97 60.30 ±$\pm 0.11$ 60.36 ±$\pm 0.12$
-M/dose_val_rx N 0.02 0.04 6.13 6.38 109.53 ±$\pm 0.20$ 105.76 ±$\pm 0.57$
-M/form_val_disp N 0.02 0.03 5.59 5.99 129.93 ±$\pm 13.27$ 82.49 ±$\pm 4.43$
+I/isopenbag C 0.01 0.01 0.00 0.00 0.00 $\pm 0.00$ 0.00 $\pm 0.00$
+I/amount N 0.03 0.08 0.00 0.05 88.26 $\pm 0.87$ 87.75 $\pm 2.87$
+I/rate N 0.04 0.13 37.49 39.54 36.47 $\pm 1.69$ 62.34 $\pm 1.11$
+I/patientweight N 0.09 - 0.00 0.00 21.23 $\pm 0.01$ 21.78 $\pm 0.02$
+I/totalamount N 0.02 0.07 12.38 13.63 6.77 $\pm 0.08$ 8.81 $\pm 0.02$
+I/originalamount N 0.03 0.12 0.17 0.10 28.00 $\pm 1.23$ 30.78 $\pm 0.34$
+I/originalrate N 0.03 0.09 5.83 4.46 41.72 $\pm 3.88$ 48.13 $\pm 2.02$
+M/drug C 0.08 - 0.00 0.00 31.00 $\pm 0.38$ 40.53 $\pm 1.01$
+M/drug_type C 0.00 0.00 0.00 0.00 0.82 $\pm 0.07$ 1.41 $\pm 0.04$
+M/prod_strength C 0.08 0.05 0.09 0.04 17.69 $\pm 0.63$ 37.79 $\pm 0.43$
+M/dose_unit_rx C 0.03 0.02 0.09 0.04 5.93 $\pm 0.11$ 13.06 $\pm 0.28$
+M/form_unit_disp C 0.03 0.04 0.09 0.05 8.16 $\pm 0.21$ 21.07 $\pm 0.45$
+M/route C 0.04 0.04 0.09 0.04 31.99 $\pm 0.49$ 40.59 $\pm 0.30$
+M/doses_per_24_hrs N 0.02 0.06 59.97 61.97 60.30 $\pm 0.11$ 60.36 $\pm 0.12$
+M/dose_val_rx N 0.02 0.04 6.13 6.38 109.53 $\pm 0.20$ 105.76 $\pm 0.57$
+M/form_val_disp N 0.02 0.03 5.59 5.99 129.93 $\pm 13.27$ 82.49 $\pm 4.43$
 Table 12: Column-wise fidelity for eICU.
@@
 Col. Name Type CDE $\downarrow$ I-CDE $\downarrow$ Real Null (%) Syn Null (%) TRTR TSTR
-L/labname C 0.04 - 0.00 0.00 52.57 ±$\pm 3.92$ 53.10 ±$\pm 12.26$
-L/labmeasurenamesystem C 0.03 0.07 5.24 4.82 0.07 ±$\pm 0.02$ 0.59 ±$\pm 0.07$
-L/labmeasurenameinterface C 0.04 0.14 6.37 5.59 54.97 ±$\pm 6.30$ 49.94 ±$\pm 0.49$
-L/labtypeid C 0.03 0.06 0.00 0.00 0.00 ±$\pm 0.00$ 0.28 ±$\pm 0.04$
-L/labresult N 0.01 0.11 0.93 0.64 14.68 ±$\pm 0.55$ 24.54 ±$\pm 0.65$
-L/labresulttext N 0.01 0.12 0.75 0.54 19.72 ±$\pm 0.46$ 23.40 ±$\pm 0.35$
-I/drugname C 0.16 - 0.00 0.00 57.38 ±$\pm 0.32$ 60.48 ±$\pm 0.20$
-I/infusionrate N 0.03 0.43 55.54 63.14 77.70 ±$\pm 0.27$ 77.95 ±$\pm 1.08$
-I/drugamount N 0.06 0.32 67.48 74.43 71.46 ±$\pm 1.57$ 69.29 ±$\pm 0.18$
-I/volumeoffluid N 0.03 0.23 67.47 74.37 14.80 ±$\pm 0.47$ 18.11 ±$\pm 0.18$
-I/patientweight N 0.03 - 90.91 95.91 12.49 ±$\pm 0.20$ 33.41 ±$\pm 0.44$
-I/drugrate N 0.03 0.36 0.46 0.55 93.12 ±$\pm 0.66$ 92.25 ±$\pm 0.25$
-M/drugname C 0.10 - 32.53 29.65 73.75 ±$\pm 0.17$ 81.10 ±$\pm 0.47$
-M/drugivadmixture C 0.02 0.03 0.00 0.00 14.49 ±$\pm 0.07$ 16.36 ±$\pm 0.01$
-M/dosage C 0.07 0.15 11.13 11.38 80.19 ±$\pm 0.08$ 81.74 ±$\pm 0.17$
-M/routeadmin C 0.07 0.11 0.00 0.01 55.23 ±$\pm 0.15$ 57.59 ±$\pm 0.09$
-M/frequency C 0.08 0.22 12.60 11.31 85.52 ±$\pm 0.19$ 88.22 ±$\pm 0.10$
-M/prn C 0.00 0.06 0.01 0.08 10.40 ±$\pm 0.12$ 11.97 ±$\pm 0.10$
+L/labname C 0.04 - 0.00 0.00 52.57 $\pm 3.92$ 53.10 $\pm 12.26$
+L/labmeasurenamesystem C 0.03 0.07 5.24 4.82 0.07 $\pm 0.02$ 0.59 $\pm 0.07$
+L/labmeasurenameinterface C 0.04 0.14 6.37 5.59 54.97 $\pm 6.30$ 49.94 $\pm 0.49$
+L/labtypeid C 0.03 0.06 0.00 0.00 0.00 $\pm 0.00$ 0.28 $\pm 0.04$
+L/labresult N 0.01 0.11 0.93 0.64 14.68 $\pm 0.55$ 24.54 $\pm 0.65$
+L/labresulttext N 0.01 0.12 0.75 0.54 19.72 $\pm 0.46$ 23.40 $\pm 0.35$
+I/drugname C 0.16 - 0.00 0.00 57.38 $\pm 0.32$ 60.48 $\pm 0.20$
+I/infusionrate N 0.03 0.43 55.54 63.14 77.70 $\pm 0.27$ 77.95 $\pm 1.08$
+I/drugamount N 0.06 0.32 67.48 74.43 71.46 $\pm 1.57$ 69.29 $\pm 0.18$
+I/volumeoffluid N 0.03 0.23 67.47 74.37 14.80 $\pm 0.47$ 18.11 $\pm 0.18$
+I/patientweight N 0.03 - 90.91 95.91 12.49 $\pm 0.20$ 33.41 $\pm 0.44$
+I/drugrate N 0.03 0.36 0.46 0.55 93.12 $\pm 0.66$ 92.25 $\pm 0.25$
+M/drugname C 0.10 - 32.53 29.65 73.75 $\pm 0.17$ 81.10 $\pm 0.47$
+M/drugivadmixture C 0.02 0.03 0.00 0.00 14.49 $\pm 0.07$ 16.36 $\pm 0.01$
+M/dosage C 0.07 0.15 11.13 11.38 80.19 $\pm 0.08$ 81.74 $\pm 0.17$
+M/routeadmin C 0.07 0.11 0.00 0.01 55.23 $\pm 0.15$ 57.59 $\pm 0.09$
+M/frequency C 0.08 0.22 12.60 11.31 85.52 $\pm 0.19$ 88.22 $\pm 0.10$
+M/prn C 0.00 0.06 0.01 0.08 10.40 $\pm 0.12$ 11.97 $\pm 0.10$
 F.2 Task-specific predictive performance
@@
 Task TRTR TSTR TRTR TSTR TRTR TSTR TRTR TSTR
-Cr 0.89 ±$\pm 0.00$ 0.88 ±$\pm 0.00$ 0.97 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.86 ±$\pm 0.00$ 0.83 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.89 ±$\pm 0.00$
-Plt 0.88 ±$\pm 0.01$ 0.87 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.84 ±$\pm 0.00$ 0.83 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.91 ±$\pm 0.00$
-WBC 0.78 ±$\pm 0.01$ 0.73 ±$\pm 0.03$ 0.91 ±$\pm 0.00$ 0.86 ±$\pm 0.00$ 0.73 ±$\pm 0.01$ 0.70 ±$\pm 0.02$ 0.87 ±$\pm 0.00$ 0.80 ±$\pm 0.00$
-HB 0.75 ±$\pm 0.01$ 0.74 ±$\pm 0.01$ 0.85 ±$\pm 0.00$ 0.82 ±$\pm 0.00$ 0.74 ±$\pm 0.00$ 0.72 ±$\pm 0.01$ 0.81 ±$\pm 0.00$ 0.78 ±$\pm 0.00$
-HCO3 0.79 ±$\pm 0.01$ 0.79 ±$\pm 0.01$ 0.90 ±$\pm 0.00$ 0.90 ±$\pm 0.00$ 0.78 ±$\pm 0.01$ 0.76 ±$\pm 0.00$ 0.88 ±$\pm 0.00$ 0.86 ±$\pm 0.00$
-Na 0.80 ±$\pm 0.01$ 0.78 ±$\pm 0.01$ 0.93 ±$\pm 0.00$ 0.92 ±$\pm 0.00$ 0.80 ±$\pm 0.01$ 0.79 ±$\pm 0.01$ 0.91 ±$\pm 0.00$ 0.90 ±$\pm 0.00$
-MgSO4 0.81 ±$\pm 0.02$ 0.81 ±$\pm 0.01$ 0.86 ±$\pm 0.00$ 0.83 ±$\pm 0.00$ 0.73 ±$\pm 0.01$ 0.72 ±$\pm 0.01$ 0.81 ±$\pm 0.01$ 0.75 ±$\pm 0.00$
-Hep 0.68 ±$\pm 0.01$ 0.65 ±$\pm 0.00$ 0.80 ±$\pm 0.00$ 0.74 ±$\pm 0.00$ 0.65 ±$\pm 0.00$ 0.63 ±$\pm 0.01$ 0.72 ±$\pm 0.00$ 0.66 ±$\pm 0.00$
-KCl 0.73 ±$\pm 0.01$ 0.70 ±$\pm 0.00$ 0.80 ±$\pm 0.00$ 0.73 ±$\pm 0.00$ 0.74 ±$\pm 0.01$ 0.66 ±$\pm 0.00$ 0.77 ±$\pm 0.00$ 0.68 ±$\pm 0.00$
-NE 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.94 ±$\pm 0.00$
-Prop 0.94 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.94 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.95 ±$\pm 0.00$ 0.96 ±$\pm 0.00$ 0.94 ±$\pm 0.00$
-Avg 0.82 ±$\pm 0.09$ 0.80 ±$\pm 0.09$ 0.90 ±$\pm 0.06$ 0.87 ±$\pm 0.08$ 0.80 ±$\pm 0.09$ 0.78 ±$\pm 0.10$ 0.87 ±$\pm 0.08$ 0.83 ±$\pm 0.10$
+Cr 0.89 $\pm 0.00$ 0.88 $\pm 0.00$ 0.97 $\pm 0.00$ 0.95 $\pm 0.00$ 0.86 $\pm 0.00$ 0.83 $\pm 0.00$ 0.94 $\pm 0.00$ 0.89 $\pm 0.00$
+Plt 0.88 $\pm 0.01$ 0.87 $\pm 0.00$ 0.96 $\pm 0.00$ 0.95 $\pm 0.00$ 0.84 $\pm 0.00$ 0.83 $\pm 0.00$ 0.94 $\pm 0.00$ 0.91 $\pm 0.00$
+WBC 0.78 $\pm 0.01$ 0.73 $\pm 0.03$ 0.91 $\pm 0.00$ 0.86 $\pm 0.00$ 0.73 $\pm 0.01$ 0.70 $\pm 0.02$ 0.87 $\pm 0.00$ 0.80 $\pm 0.00$
+HB 0.75 $\pm 0.01$ 0.74 $\pm 0.01$ 0.85 $\pm 0.00$ 0.82 $\pm 0.00$ 0.74 $\pm 0.00$ 0.72 $\pm 0.01$ 0.81 $\pm 0.00$ 0.78 $\pm 0.00$
+HCO3 0.79 $\pm 0.01$ 0.79 $\pm 0.01$ 0.90 $\pm 0.00$ 0.90 $\pm 0.00$ 0.78 $\pm 0.01$ 0.76 $\pm 0.00$ 0.88 $\pm 0.00$ 0.86 $\pm 0.00$
+Na 0.80 $\pm 0.01$ 0.78 $\pm 0.01$ 0.93 $\pm 0.00$ 0.92 $\pm 0.00$ 0.80 $\pm 0.01$ 0.79 $\pm 0.01$ 0.91 $\pm 0.00$ 0.90 $\pm 0.00$
+MgSO4 0.81 $\pm 0.02$ 0.81 $\pm 0.01$ 0.86 $\pm 0.00$ 0.83 $\pm 0.00$ 0.73 $\pm 0.01$ 0.72 $\pm 0.01$ 0.81 $\pm 0.01$ 0.75 $\pm 0.00$
+Hep 0.68 $\pm 0.01$ 0.65 $\pm 0.00$ 0.80 $\pm 0.00$ 0.74 $\pm 0.00$ 0.65 $\pm 0.00$ 0.63 $\pm 0.01$ 0.72 $\pm 0.00$ 0.66 $\pm 0.00$
+KCl 0.73 $\pm 0.01$ 0.70 $\pm 0.00$ 0.80 $\pm 0.00$ 0.73 $\pm 0.00$ 0.74 $\pm 0.01$ 0.66 $\pm 0.00$ 0.77 $\pm 0.00$ 0.68 $\pm 0.00$
+NE 0.96 $\pm 0.00$ 0.96 $\pm 0.00$ 0.96 $\pm 0.00$ 0.96 $\pm 0.00$ 0.96 $\pm 0.00$ 0.95 $\pm 0.00$ 0.95 $\pm 0.00$ 0.94 $\pm 0.00$
+Prop 0.94 $\pm 0.00$ 0.94 $\pm 0.00$ 0.95 $\pm 0.00$ 0.94 $\pm 0.00$ 0.96 $\pm 0.00$ 0.95 $\pm 0.00$ 0.96 $\pm 0.00$ 0.94 $\pm 0.00$
+Avg 0.82 $\pm 0.09$ 0.80 $\pm 0.09$ 0.90 $\pm 0.06$ 0.87 $\pm 0.08$ 0.80 $\pm 0.09$ 0.78 $\pm 0.10$ 0.87 $\pm 0.08$ 0.83 $\pm 0.10$
 Table 14: Comparison of VQ-VAE and RQ-VAE on MIMIC-IV and eICU, reporting Column-wise Density Estimation (CDE), Pairwise Column Correlation (PCC), Time Gap, and Event Count metrics.
@@
 Model CDE $\downarrow$ PCC $\downarrow$ Time Gap $\downarrow$ # Events $\downarrow$ Utility $\uparrow$
-Real 0.83±$\pm 0.00$
-SDV 0.05 0.21 0.46 0.07 0.63±$\pm 0.09$
-RC-TGAN 0.55 0.20 0.34 0.06 0.57±$\pm 0.06$
-ClavaDDPM 0.16 0.11 0.35 0.06 0.64±$\pm 0.06$
-RawMed 0.08 0.02 0.02 0.02 0.79±$\pm 0.05$
+[…]
+[…]
+[…]
+[…]
+[…]
 F.3 ED dataset generalizability
@@
   • Heart rate (beats per minute, bpm): 1 if the maximum value is $\geq 120$ bpm (indicating tachycardia, above the normal adult range of 60–100 bpm), 0 otherwise.
-  • Respiratory rate (respirations per minute, rpm): 1 if the maximum value is $>24$ rpm (indicating tachypnea, beyond the normal adult range of 12–20 rpm), 0 otherwise.
+  • Respiratory rate (respirations per minute, rpm): 1 if the maximum value is >24>24 rpm (indicating tachypnea, beyond the normal adult range of 12–20 rpm), 0 otherwise.
   • Morphine administration: 1 if administered, 0 if not.
@@
 Despite differing clinical environments, RawMed’s performance closely matches real data utility and surpasses baselines, indicating potential for generalization to other datasets.
+F.4 Conditional generation
+We conducted experiments on conditional generation by incorporating static features into the Temporal Modeling framework (Section 3.3).
+Specifically, we integrated age, gender, and admission diagnosis (or type) as static features.
+During training, these features were prepended to the sequences, enabling the model to learn the combined sequence auto-regressively.
+For generation, the model utilized real static features as conditions to produce the complete sequence.
+To further investigate conditional generation, we explored the inclusion of initial medical context, providing the model with real static features and approximately one-fourth of the initial real-data events from the sequence, tasking it with generating the remaining three-fourths.
+To assess whether the generated sequences effectively captured the relationships with static features, we trained three classifiers to predict each age, gender, admission diagnosis or type from the generated sequences, excluding the static feature portion itself.
+For this classification, age was binned, and for the admission diagnosis, only four specific categories from the ’Non-operative Organ Systems’ hierarchy (which account for the majority of diagnoses) were selected, namely cardiovascular, neurologic, respiratory, and gastrointestinal.
+Results
+The performance of our conditional generation approach on the MIMIC-IV and eICU datasets is presented in Table 17.
+For the MIMIC-IV dataset, the synthetic data demonstrates performance comparable to real data, indicating that our model effectively captures the relationship between static features and sequential data.
+Notably, conditioning on initial events enhances performance, with the AUROC for age prediction increasing from 0.75 to 0.77.
+For gender prediction, both real and synthetic data achieved an AUROC of 1.00.
+This is attributable to the inclusion of all columns from the labevents table in our synthetic data generation, which unintentionally incorporates gender-specific reference ranges in lab values (e.g., creatinine, hemoglobin) within the ref_range_lower and ref_range_upper columns, serving as direct gender indicators.
+On the eICU dataset, similar trends were observed, with synthetic data approaching the performance of real data, particularly when leveraging initial medical context.
+The AUROC values improved when conditioning on initial events, rising from 0.65 to 0.69 for age, 0.52 to 0.59 for gender, and 0.86 to 0.88 for admission diagnosis, compared to static features alone.
+These results underscore the flexibility of our framework in supporting conditional generation with richer feature sets, enhancing synthetic data quality.
+Although gender prediction on eICU exhibited a slight performance drop compared to real data, we anticipate improvements through further methodological refinements and hyperparameter optimization.
+The results from both datasets validates the flexibility and potential of our approach for conditional generation.
 (a) MIMIC-IV / lab
@@
 Figure 4: Cumulative Distribution Functions (CDFs) of real and synthetic data for MIMIC-IV and eICU datasets, showing Time Gap (with and without simultaneous events), Absolute Time from admission, and Event Counts per patient.
+Table 17: Performance (AUROC) of conditional generation models on MIMIC-IV and eICU datasets. Adm. Type refers to Admission Type, and Adm. Diag. refers to Admission Diagnosis.
+MIMIC-IV eICU
+Model Age Gender Adm. Type Age Gender Adm. Diag.
+Real 0.80 1.00 0.92 0.72 0.67 0.89
+RawMed (static only) 0.75 1.00 0.89 0.65 0.52 0.86
+RawMed (static+1/4 initial events) 0.77 1.00 0.89 0.69 0.59 0.88